### v1

# CFG++: Manifold-constrained Classifier Free Guidance for Diffusion Models

Abstract
Classifier-free guidance (CFG) is a fundamental tool in modern diffusion models for text-guided generation. Although effective, CFG has notable drawbacks. For instance, DDIM with CFG lacks invertibility, complicating image editing; furthermore, high guidance scales, essential for high-quality outputs, frequently result in issues like mode collapse. Contrary to the widespread belief that these are inherent limitations of diffusion models, this paper reveals that the problems actually stem from the off-manifold phenomenon associated with CFG, rather than the diffusion models themselves. More specifically, inspired by the recent advancements of diffusion model-based inverse problem solvers (DIS), we reformulate text-guidance as an inverse problem with a text-conditioned score matching loss, and develop CFG++, a novel approach that tackles the off-manifold challenges inherent in traditional CFG. CFG++ features a surprisingly simple fix to CFG, yet it offers significant improvements, including better sample quality for text-to-image generation, invertibility, smaller guidance scales, reduced mode collapse, etc. Furthermore, CFG++ enables seamless interpolation between unconditional and conditional sampling at lower guidance scales, consistently outperforming traditional CFG at all scales. Experimental results confirm that our method significantly enhances performance in text-to-image generation, DDIM inversion, editing, and solving inverse problems, suggesting a wide-ranging impact and potential applications in various fields that utilize text guidance. Project Page: https://cfgpp-diffusion.github.io/.
Figure 1: (Top) Comparison of T2I results by SDXL for the prompt "kayak in the water, optical color, aerial view, rainbow". The CFG-guided image has a broken oar, an artifact significantly reduced in the CFG++ version. (Middle) DDIM Inversion results under CFG show noticeable artifacts at various CFG scales, which are significantly reduced by CFG++. (Bottom) The evolution of denoised estimates differs between CFG and CFG++. CFG exhibits sudden shifts and intense color saturation early in reverse diffusion, while CFG++ transitions smoothly from low to high-resolution details.
1 Introduction
Classifier-free guidance (CFG) (Ho & Salimans, 2021) forms the key basis of modern text-guided generation with diffusion models (Dhariwal & Nichol, 2021; Rombach et al., 2022). Nowadays, it is common practice to train a diffusion model with large-scale paired text-image data (Schuhmann et al., 2022), so that sampling (i.e. generating) a signal (e.g. image, video) from a diffusion model can either be done unconditionally from $p_{\theta}({\boldsymbol{x}}|\varnothing)\equiv p_{\theta}({\boldsymbol{x}})$, or conditionally from $p_{\theta}({\boldsymbol{x}}|{\boldsymbol{c}})$, where ${\boldsymbol{c}}$ is the text conditioning. Once trained, it seems natural that one would acquire samples from the conditional distribution by simply solving the probability-flow ODE or SDE sampling (Song et al., 2021b; a; Karras et al., 2022) with the conditional score function. In practice, however, it is observed that the conditioning signal is insufficient when used naively. To emphasize the guidance, one uses the guidance scale $\omega>1$, where the direction can be defined by the direction from the unconditional score to the conditional score (Ho & Salimans, 2021).
In modern text-to-image (T2I) diffusion models, the guidance scale $\omega$ is typically set within the range of $[5.0,30]$, referred to as the moderately high range of CFG guidance  (Chen et al., 2024b; Podell et al., 2023). The insufficiency in guidance also holds for classifier guidance (Dhariwal & Nichol, 2021; Song et al., 2021b) so that a scale of 10 was used. While using a high guidance scale yields higher-quality images with better alignment to the condition, it is also prone to mode collapse, reduces sample diversity, and yields an inevitable accumulation of errors during the sampling process. One example is DDIM inversion (Dhariwal & Nichol, 2021), a pivotal technique for controllable synthesis and editing (Mokady et al., 2023), where running the inversion process with $\omega>1.0$ leads to significant compromise in the reconstruction performance (Mokady et al., 2023; Wallace et al., 2023). Another extreme example would be score distillation sampling (SDS) (Poole et al., 2022), where the guidance scale in the order of a few hundred is chosen. Using such a high guidance scale leads to better asset quality to some extent, but induces blurry and saturated results. Several research efforts have been made to mitigate this downside by exploring methods where using a smaller guidance scale suffices (Wang et al., 2024; Liang et al., 2023). Although recent progress in SDS-type methods has reduced the necessary guidance scale to a range that is similar to those of ancestral samplers, using a moderately large $\omega$ is considered an inevitable choice.
In this work, we aim to give an answer to this conundrum by revisiting the geometric view of diffusion models. In particular, inspired by the recent advances in diffusion-based inverse problem solvers (DIS) (Kadkhodaie & Simoncelli, 2021; Chung et al., 2023a; Song et al., 2023a; Kim et al., 2024; Chung et al., 2024), we reformulate the text-guidance as an inverse problem with a text-conditioned score matching loss and derive reverse diffusion sampling strategy by utilizing the decomposed diffusion sampling (DDS) (Chung et al., 2024). This results in a surprisingly simple fix of CFG to the sampling process without any computational overhead. The resulting process, which we call CFG++, works with a small guidance scale, typically $\lambda\in[0.0,1.0]$, that smoothly interpolates between unconditional and conditional sampling, with $\lambda=1.0$ having a similar effect as using CFG sampling with $\omega\sim 12.5$ at 50 neural function evaluation (NFE). Furthermore, DDIM inversion with CFG++ is invertible up to the discretization error, simplifying image editing. Comparing CFG++ against CFG shows that we achieve consistently better sample quality for text-to-image (T2I) generation, significantly better DDIM inversion capabilities that lead to enhanced reconstruction and editing, and enabling the incorporation of CFG guidance to diffusion inverse solvers (DIS) (Chung et al., 2023a). While the applications of CFG++ that we show in this work are limited, we believe that our work will have a broad impact that can be applied to all applications that leverage text guidance through the traditional CFG.
Algorithm 1 Reverse Diffusion with CFG
1:${\boldsymbol{x}}_{T}\sim{\mathcal{N}}(0,{\mathbf{I}}_{d}),0\leq{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\omega}\in\mathbb{R}$
2:for $i=T$ to $1$ do
3:     ${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})={\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})+{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\omega}[{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})]$
4:     ${\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})\leftarrow({\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t}))/\sqrt{\bar{\alpha}_{t}}$
5:     ${\boldsymbol{x}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$
6:end for
7:return ${\boldsymbol{x}}_{0}$
Algorithm 2 Reverse Diffusion with CFG++
${\boldsymbol{x}}_{T}\sim{\mathcal{N}}(0,{\mathbf{I}}_{d}),{\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\lambda}\in[0,1]$
2:for $i=T$ to $1$ do
3:     ${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})={\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})+{\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\lambda}[{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})]$
4:     ${\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})\leftarrow({\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t}))/\sqrt{\bar{\alpha}_{t}}$
5:     ${\boldsymbol{x}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}{\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})$
6:end for
7:return ${\boldsymbol{x}}_{0}$
Figure 2: Comparison between reverse diffusion process by CFG and CFG++. CFG++ proposes a simple but surprisingly effective fix: using ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})$ instead of ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$ in updating ${\boldsymbol{x}}_{t-1}$.
2 Background
Diffusion models. Diffusion models (Ho et al., 2020; Song et al., 2021b; Karras et al., 2022) are generative models designed to learn the reversal of a forward noising process. This process starts with an initial distribution $p_{0}({\boldsymbol{x}})$ where ${\boldsymbol{x}}\in{\mathbb{R}}^{n}$, and progresses towards the standard Gaussian distribution $p_{T}({\boldsymbol{x}})\approx\mathcal{N}(\bm{0},{\bm{I}})$, utilizing forward Gaussian perturbation kernels. Sampling from the data distribution can be performed by solving either the reverse stochastic differential equation (SDE) or the equivalent probability-flow ordinary differential equation (PF-ODE) (Song et al., 2021b). For example, under the choice $p({\boldsymbol{x}}_{t}|{\boldsymbol{x}}_{0})=\mathcal{N}({\boldsymbol{x}}_{0},t^{2}{\bm{I}})$, the generative PF-ODE reads
$d{\boldsymbol{x}}_{t}=-t\nabla_{{\boldsymbol{x}}_{t}}\log p({\boldsymbol{x}}_{t})\,dt=\frac{{\boldsymbol{x}}_{t}-\mathbb{E}[{\boldsymbol{x}}_{0}|{\boldsymbol{x}}_{t}]}{t}\,dt,\,\,{\boldsymbol{x}}_{T}\sim p_{T}({\boldsymbol{x}}_{T}),$ (1)
where Tweedie’s formula (Efron, 2011) ${\mathbb{E}}[{\boldsymbol{x}}_{0}|{\boldsymbol{x}}_{t}]={\boldsymbol{x}}_{t}+t^{2}\nabla_{{\boldsymbol{x}}_{t}}\log p({\boldsymbol{x}}_{t})$ is applied to achieve the second equality. Thus, the score function $\nabla_{{\boldsymbol{x}}_{t}}\log p({\boldsymbol{x}}_{t})$ plays the key role in diffusion models. One can approximate an equivalent parameterization of the score function via denoising score matching (Hyvärinen & Dayan, 2005) or epsilon-matching (Ho et al., 2020), etc. When aiming for a text-conditional diffusion model that can condition on arbitrary ${\boldsymbol{c}}$, one can extend epsilon matching to a conditional one as:
$\min_{\theta}{\mathbb{E}}_{{\boldsymbol{x}}_{t}\sim p({\boldsymbol{x}}_{t}|{\boldsymbol{x}}_{0}),\{{\boldsymbol{x}}_{0},{\boldsymbol{c}}\}\sim p({\boldsymbol{x}}_{0},{\boldsymbol{c}}),{\boldsymbol{\epsilon}}\sim{\mathcal{N}}(0,{\bm{I}})}\left[\|{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},{\boldsymbol{c}})-{\boldsymbol{\epsilon}}\|_{2}^{2}\right],$ (2)
where the condition is dropped with a certain probability (Ho & Salimans, 2021) so that null conditioning with ${\boldsymbol{c}}=\varnothing$ is possible. The neural network architecture of ${\boldsymbol{\epsilon}}_{\theta}$ is designed so that the condition ${\boldsymbol{c}}$ can modulate the output through cross attention (Rombach et al., 2022). For simplicity in notation throughout the paper, we define ${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}:={\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},{\boldsymbol{c}})$ and ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}:={\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},\varnothing)$ by dropping $\theta$ and ${\boldsymbol{x}}_{t}$.
Extending the result of the Tweedie’s formula to the unconditional case under the variance preserving (VP) framework of DDPMs (Ho et al., 2020), we have
${\mathbb{E}}[{\boldsymbol{x}}_{0}|{\boldsymbol{x}}_{t},\varnothing]=\hat{\boldsymbol{x}}_{\varnothing}({\boldsymbol{x}}_{t}):=({\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t}))/\sqrt{\bar{\alpha}_{t}}.$ (3)
Leveraging (3), it is common to use DDIM sampling (Song et al., 2021b) to solve the conditional probability-flow ODE (PF-ODE) of the generative process. Specifically, a single iteration reads
${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}$ $=({\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}})/\sqrt{\bar{\alpha}_{t}}$ (4)
${\boldsymbol{x}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}},$ (5)
where ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}:=\hat{\boldsymbol{x}}_{\varnothing}({\boldsymbol{x}}_{t})={\mathbb{E}}[{\boldsymbol{x}}_{0}|{\boldsymbol{x}}_{t},\varnothing]$ is the denosed signal by Tweedie’s formula, and (5) corresponds to the renoising step. This is repeated for $t=T,T-1,\dots,1$.
For modern diffusion models, it is common to train a diffusion model in the latent space (Rombach et al., 2022) with the latent variable ${\boldsymbol{z}}$. While most of our experiments are performed with latent diffusion models (LDM), as our framework holds both for pixel- and latent- diffusion models, we will simply use the notation ${\boldsymbol{x}}$ regardless of the choice.
Classifier free guidance. For a conditional diffusion, ho2021classifierfree considered the sharpened posterior distribution $p^{\omega}({\boldsymbol{x}}|{\boldsymbol{c}})\propto p({\boldsymbol{x}})p({\boldsymbol{c}}|{\boldsymbol{x}})^{\omega}$. Using Bayes rule for some timestep $t$,
$\nabla_{\boldsymbol{x}}\log p^{\omega}({\boldsymbol{x}}_{t}|{\boldsymbol{c}})=\nabla_{{\boldsymbol{x}}_{t}}\log p({\boldsymbol{x}}_{t})+\omega(\nabla_{{\boldsymbol{x}}_{t}}\log p({\boldsymbol{x}}_{t}|{\boldsymbol{c}})-\nabla_{{\boldsymbol{x}}_{t}}\log p({\boldsymbol{x}}_{t}))$ (6)
Parametrizing the score function with ${\boldsymbol{\epsilon}}_{\theta}$, as in DDPM (ho2020denoising), we have
${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t}):={\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})+\omega[{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})]$ (7)
where we introduce a compact notation ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}$ that guides the sampling from the sharpened posterior. When sampling with CFG guidance with DDIM sampling, one replaces ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}$ with ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}$ for both the Tweedie estimate (4) and the subsequent update step (5), leading to Algorithm 1.
Diffusion model-based inverse problem solvers. Diffusion model-based inverse problem solvers (DIS) aim to perform posterior sampling from an unconditional diffusion model (kadkhodaie2021stochastic; chung2023diffusion; song2023pseudoinverseguided; kim2024dreamsampler). Specifically, for a given loss function $\ell({\boldsymbol{x}})$ which often stems from the likelihood for measurement consistency [^1] For example, for a given measurement model ${\boldsymbol{y}}=A{\boldsymbol{x}}+{\boldsymbol{w}}$ with ${\boldsymbol{w}}\sim\mathcal{N}({\boldsymbol{x}}_{0},\sigma^{2}{\bm{I}})$, the loss function is the data consistency (DC) loss given by $\ell({\boldsymbol{x}})=\|{\boldsymbol{y}}-A{\boldsymbol{x}}\|^{2}/2\sigma^{2}$. , the goal of DIS is to address the following optimization problem
$\min_{{\boldsymbol{x}}\in{\mathcal{M}}}\ell({\boldsymbol{x}})$ (8)
where ${\mathcal{M}}$ represents the clean data manifold sampled from unconditional distribution $p_{0}({\boldsymbol{x}})$. Consequently, it is essential to navigate in a way that minimizes cost while also identifying the correct clean manifold.
Recently, chung2023diffusion proposed a general technique called diffusion posterior sampling (DPS), where the updated estimate from the noisy sample ${\boldsymbol{x}}_{t}\in{\mathcal{M}}_{t}$ is constrained to stay on the same noisy manifold ${\mathcal{M}}_{t}$. This is achieved by computing the manifold constrained gradient (MCG) (chung2022improving) on a noisy sample ${\boldsymbol{x}}_{t}\in{\mathcal{M}}_{t}$ as $\nabla^{mcg}_{{\boldsymbol{x}}_{t}}\ell({\boldsymbol{x}}_{t}):=\nabla_{{\boldsymbol{x}}_{t}}\ell(\hat{\boldsymbol{x}}_{t})$, where $\hat{\boldsymbol{x}}_{t}$ is the denoised sample in (3) through Tweedie’s formula. The resulting algorithm can be stated as follows:
${\boldsymbol{x}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\left({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}-\gamma_{t}\nabla_{{\boldsymbol{x}}_{t}}\ell({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})\right)+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}},$ (9)
where $\gamma_{t}>0$ denotes the step size. Under the linear manifold assumption (chung2022improving; chung2023diffusion), this allows precise transition to ${\mathcal{M}}_{t-1}$. Therefore, by performing (9) from $t=T$ to $t=0$, we can solve the optimization problem (8) with ${\boldsymbol{x}}_{0}\in{\mathcal{M}}$. Unfortunately, the computation of MCG requires computationally expensive backpropagation and is often unstable  (poole2022dreamfusion; du2023reduce).
In another work, chung2024decomposed shows that (9) can be equivalently represented as
${\boldsymbol{x}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}P_{\mathcal{M}}\left({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}-\gamma_{t}\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})\right)+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}},$ (10)
where $P_{\mathcal{M}}$ denotes the projection on the clean manifold ${\mathcal{M}}$. This implies that the data consistency update step becomes a standard gradient step, followed by the projection on the clean manifold. Importantly, under the commonly used linear manifold assumption in DPS, the one step update by ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}-\gamma_{t}\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})$ or multi-step update employing the conjugate gradient (CG) method are guaranteed to remain within a linear subspace, thus obviating the need for explicit computation of $P_{\mathcal{M}}$, i.e.
${\boldsymbol{x}}_{t-1}$ $\simeq\sqrt{\bar{\alpha}_{t-1}}\left({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}-\gamma_{t}\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})\right)+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}.$ (11)
This method, often referred to as the decomposed diffusion sampling (DDS), bypasses the computation of the score Jacobian, similar to poole2022dreamfusion, making it stable and suitable for large scale medical imaging inverse problems (chung2024decomposed). In the following, we leverage the insight from DDS to propose an algorithm to improve upon the CFG algorithm.
Figure 3: Off-manifold phenomenon of CFG arise from: (a) the typical CFG scale $\omega>1.0$ which leads to extrapolation and deviation from the piecewise linear data manifold, and (b) CFG’s renoising process, which introduces a nonzero offset $\Delta^{\omega}$ from the correct manifold. CFG++ effectively mitigates all these artifacts.
3 CFG++ : Manifold-Constrained CFG
3.1 Derivation of CFG++
Instead of uncritically adopting the sharpened posterior distribution $p^{\omega}({\boldsymbol{x}}|{\boldsymbol{c}})\propto p({\boldsymbol{x}})p({\boldsymbol{c}}|{\boldsymbol{x}})^{\omega}$ as introduced by ho2021classifierfree and further explored in the Bayesian CFG formulation, we adopt a fundamentally different strategy by reformulating text-guidance as an optimization problem. Specifically, our focus is on identifying a loss function $\ell({\boldsymbol{x}})$ in (8) such that, when minimized under the condition set by the text, enables the reverse diffusion process to progressively generate samples that increasingly satisfy the text condition.
One of the most significant contributions of this paper is to reveal that the text-conditioned score matching loss or score distillation sampling (SDS) loss is also ideally suited for our purpose. Specifically, we are interested in solving the following inverse problem through diffusion models:
$\min_{{\boldsymbol{x}}\in{\mathcal{M}}}\ell_{sds}(x),\quad\ell_{sds}({\boldsymbol{x}}):=\|{\boldsymbol{\epsilon}}_{\theta}(\sqrt{\bar{\alpha}_{t}}{\boldsymbol{x}}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}},{\boldsymbol{c}})-{\boldsymbol{\epsilon}}\|_{2}^{2}$ (12)
This implies that our goal is to identify solutions on the clean manifold ${\mathcal{M}}$ that optimally align with the text condition ${\boldsymbol{c}}$.
Figure 4: Enhanced T2I results by SDXL ($\omega=9.0,\lambda=0.8$) with CFG++. Under CFG, the lion cub is not visible (top-left), the dog appears with two tails (top-right), the goggles have an unusual shape (bottom-left), and the tree trunk is folded (bottom-right). These artifacts are absent in those produced by CFG++.
To avoid the Jacobian computation, in this paper we attempt to solve (12) through decomposed diffusion sampling (DDS). The resulting sampling process from reverse diffusion is then given by
${\boldsymbol{x}}_{t-1}$ $=$ $\sqrt{\bar{\alpha}_{t-1}}\left({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}-\gamma_{t}\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell_{sds}({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})\right)+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}$ (13)
which is analogous to (11). Consequently, the remaining technical issue is to compute $\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell_{sds}({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})$. To answer this, by using the forward diffusion to generate ${\boldsymbol{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\boldsymbol{x}}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}$ from the clean image ${\boldsymbol{x}}\in{\mathcal{M}}$, we convert the objective function in (12) into the following equivalent loss:
$\ell_{sds}({\boldsymbol{x}})$ $=\|{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},{\boldsymbol{c}})\|^{2}=\left\|\frac{{\boldsymbol{x}}_{t}-\sqrt{\bar{\alpha}_{t}}{\boldsymbol{x}}}{\sqrt{1-\bar{\alpha}_{t}}}-\frac{{\boldsymbol{x}}_{t}-\sqrt{\bar{\alpha}_{t}}{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}}{\sqrt{1-\bar{\alpha}_{t}}}\right\|^{2}$ (14)
$=\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}\|{\boldsymbol{x}}-{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}\|^{2}$ (15)
where we use the conditional Tweedie formula derived by replacing ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}$ in (4) by the conditioned score ${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}$. Similar observation has been made in DreamSampler (kim2024dreamsampler) and HiFA (zhu2023hifa). Consequently, $\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell_{sds}({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})$ can be readily computable and the resulting DDS sampling process can be simplified as
${\boldsymbol{x}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\left({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}+\lambda({\hat{\boldsymbol{x}}_{\boldsymbol{c}}}-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})\right)+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}$ (16)
where $\lambda:=\frac{2{\bar{\alpha}_{t}}}{{1-\bar{\alpha}_{t}}}\gamma_{t}$. Although concise, (16) can be further simplified to a similar form of the traditional CFG as described in Algorithm 1. Specifically, note that
${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}+\lambda({\hat{\boldsymbol{x}}_{\boldsymbol{c}}}-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})=({\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t}))/\sqrt{\bar{\alpha}_{t}}$ (17)
where
${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t}):={\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})+{\lambda}[{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})]$ (18)
Therefore, (16) can be equivalently represented by:
${\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})$ $=({\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t}))/\sqrt{\bar{\alpha}_{t}}$ (19)
${\boldsymbol{x}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}{\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})$ (20)
which is summarized in Algorithm 2.
By examining Algorithm 1 and Algorithm 2, we observe a notable similarity between CFG and CFG++. However, a crucial difference exists. Specifically, in CFG++, the renoising process after applying Tweedie’s formula should utilize the unconditional noise ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})$ instead of ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$. This surprisingly simple fix to the original CFG algorithm leads to smoother trajectory of generation, (Fig. 1 bottom) and generation with superior quality (Fig. 1 top). Moreover, CFG++ can also be easily extended to significantly improve DDIM inversion, which is known to be prone to failure by standard CFG (mokady2023null). See Fig. 1 (middle) for representative results and Sec. 4.2 for further discussions. In the following section, we present a theoretical analysis of CFG++ to investigate the origin of this improvement.
3.2 Geometry of CFG++
Mitigating off-manifold phenomenon. In Fig. 1 (bottom), we illustrate the evolution of the posterior mean through Tweedie’s formula during the reverse diffusion process. Notably, in the early phases of reverse diffusion sampling under CFG, there is a sudden shift in the image and intense color saturation, followed by only minor changes until later stages. This abrupt transition in the posterior mean is somewhat unexpected, as the reverse diffusion process is designed to progressively refine resolution, moving from a noisier to a clearer manifold. Similar phenomena are depicted in Fig. 13 in Appendix. We refer to this as the off-manifold phenomenon, which will be clarified shortly. Conversely, under CFG++, the posterior mean evolves gradually, transitioning smoothly from a low-resolution image to high-resolution details. This indicates that CFG++ is less prone to the off-manifold phenomenon observed with CFG. In the following, we provide a theoretical analysis of the origin of the off-manifold phenomenon and explain why CFG++ can mitigate this issue.
By examining (17) and (19) as well as Algorithm 1, we can easily see that the Tweedie’s formula for the CFG++ and CFG can be equivalently represented by
${\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})=(1-\lambda){\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\lambda{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t}),\quad{\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})=(1-\omega){\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\omega{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})$ (21)
This implies that $\lambda,\omega\in[0,1]$ faciliates an interpolation between the unconditional and the conditional posterior means. However, given the typical CFG scale $\omega>1.0$, we observe that CFG extrapolates beyond the unconditional and the conditional posterior means established by Tweedie’s formula. Consequently, under the usual assumption that the clean manifold to be piecewise linear around the posterior mean derived via Tweedie’s formula, the conditional posterior mean estimates from traditional CFG, obtained with a guidance scale outside the range of $[0,1]$, can extend beyond the piecewise linear manifold. This may lead to the estimates potentially ‘falling off’ the data manifold, as depicted by an orange arrow pointing downwards in Fig. 3(a). We conjecture that this is one of the primary contributors to the off-manifold phenomenon. Thus, we select $\lambda\in[0,1]$ as the guidance scale for CFG++ to ensure it remains an interpolation between the unconditional and conditional Tweedie, thus preventing it from ‘falling off’ the clean data manifold.
An additional source of the off-manifold phenomenon in CFG occurs during the transition from the clean manifold ${\mathcal{M}}$ to the subsequent noisy manifold ${\mathcal{M}}_{t-1}$. Specifically, even from the same point in the clean manifold, the use of ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$ for this transition introduces a nonzero offset $\Delta^{\omega}:=\omega\sqrt{1-\bar{\alpha}_{t-1}}({\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t}))$ from ${\mathcal{M}}_{t-1}$, as illustrated in Fig. 3 (b). This leads to the invalid lifts from ${\mathcal{M}}$ to ${\mathcal{M}}_{t-1}$. On the other hand, we note that the renoising process of CFG++ follows the standard DDIM, resulting in the accurate transition to ${\mathcal{M}}_{t-1}$.
Figure 5: The normalized text-conditioned score matching loss throughout the reverse diffusion sampling process for both CFG and CFG++.
Text-Image alignment. In Fig. 1 (top), we display images generated by SDXL using the prompt “Kayak in the water, optical color, areal view, rainbow". The generated image on the left is guided by the traditional CFG, while the one on the right is produced with our CFG++. Although both images look impressive and appear to align well with the given text prompt, a closer examination reveals a significant difference. For instance, one of the oars in the CFG-guided image is broken in the middle, which shows a misalignment with the text prompt. This type of artifact is substantially reduced in the CFG++ version, where the oars are depicted intact. The examples provided by SDXL in Fig. 4 further demonstrate the enhanced text-to-image alignment achieved with CFG++. For instance, the lion cub is not visible in Fig. 4(top-left), the dog appears with two tails in Fig. 4(top-right), the goggles have an unusual shape in Fig. 4(bottom-left), and the tree trunk is folded in Fig. 4(bottom-right). These artifacts, observed in the images generated by CFG, are absent in those produced by CFG++.
In fact, this enhanced alignment capability is a natural consequence of CFG++, which directly minimizes the text-conditioned score-matching loss as shown in (12). In contrast, CFG indirectly seeks text alignment through the sharpened posterior distribution $p^{\omega}({\boldsymbol{x}}|{\boldsymbol{c}})\propto p({\boldsymbol{x}})p({\boldsymbol{c}}|{\boldsymbol{x}})^{\omega}$. Therefore, CFG++ inherently outperform CFG in terms of text alignment due to its fundamental design principle. For example, Fig. 5 displays the normalized text-conditioned score matching loss, represented as $(1-\bar{\alpha}_{t})\|{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},{\boldsymbol{c}})\|^{2}/\bar{\alpha}_{t}=\|{\boldsymbol{x}}-{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}\|^{2}$, throughout the reverse diffusion sampling process for both CFG and CFG++. The loss plot associated with CFG shows fluctuations and maintains a noticeable gap compared to CFG++ even after the completion of the reverse diffusion process. In fact, the fluctuation associated with CFG is also related to the off-manifold issue, and from Fig. 5 we can easily see that the off-manifold phenomenon is more dominant at early stage of reverse diffusion sampling. Conversely, the loss trajectory for CFG++ demonstrates a much smoother variation, particularly during the early stages of reverse diffusion.
DDIM inversion. As discussed in (song2020denoising), the denoising process for unconditional DDIM is approximately invertible, meaning that ${\boldsymbol{x}}_{t}$ can generally be recovered from ${\boldsymbol{x}}_{t-1}$ Specifically, from (4) and (5), we have the following approximate inversion formula for unconditional DDIM:
${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})$ $=({{\boldsymbol{x}}_{t-1}-\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})})/{\sqrt{\bar{\alpha}_{t-1}}}$
$\simeq({{\boldsymbol{x}}_{t-1}-\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1})})/{\sqrt{\bar{\alpha}_{t-1}}}$ (22)
${\boldsymbol{x}}_{t}$ $=\sqrt{\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})$
$\simeq\sqrt{\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1})$ (23)
where the approximation arises from ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})\simeq{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1})$. A similar inversion procedure has been employed for conditional DDIM inversion under CFG By assuming ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})\simeq{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t-1})$, and replacing ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}$ by ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}$ as detailed in Algorithm 3.
On the other hand, by examining (19) and (20), we can obtain the following approximate DDIM inversion formula under CFG++:
${\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})$ $\simeq({{\boldsymbol{x}}_{t-1}-\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1})})/{\sqrt{\bar{\alpha}_{t-1}}}$ (24)
${\boldsymbol{x}}_{t}$ $\simeq\sqrt{\bar{\alpha}_{t}}{\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t}}{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t-1})$ (25)
where we apply the approximation ${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})\simeq{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t-1})$ alongside the usual assumption ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})\simeq{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1})$. This formulation underpins the CFG++ guided DDIM inversion algorithm, presented in Algorithm 4.
Algorithm 3 DDIM Inversion with CFG
1:${\boldsymbol{x}}_{0},0\leq{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\omega}\in\mathbb{R}$
2:for $i=0$ to $T-1$ do
3:     ${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})={\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})+{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\omega}[{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})]$
4:     ${\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})\leftarrow({\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t}))/\sqrt{\bar{\alpha}_{t}}$
5:     ${\boldsymbol{x}}_{t+1}=\sqrt{\bar{\alpha}_{t+1}}{\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t+1}}{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$
6:end for
7:return ${\boldsymbol{x}}_{T}$
Algorithm 4 DDIM Inversion with CFG++
${\boldsymbol{x}}_{0},{\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\lambda}\in[0,1]$
2:for $i=0$ to $T-1$ do
3:     ${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})={\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})+{\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\lambda}[{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})]$
4:     ${\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})\leftarrow({\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t}))/\sqrt{\bar{\alpha}_{t}}$
5:     ${\boldsymbol{x}}_{t+1}=\sqrt{\bar{\alpha}_{t+1}}{\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t+1}}{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})$
6:end for
7:return ${\boldsymbol{x}}_{T}$
Figure 6: Comparison between DDIM inversion. CFG++ proposes a simple yet effective fix: using ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})$ instead of ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$ in Tweedie’s denoising step.
In practice, for small time step size the error ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})\simeq{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1})$ is relatively small, so the unconditional DDIM inversion by (22) and (23) lead to relatively insignificant errors. Unfortunately, the corresponding inversion from conditional diffusion under CFG is quite distorted as noted in (mokady2023null; wallace2023edict). In fact, this distortion is originated from the inaccuracy of the approximation ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})\simeq{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t-1})$. More specifically, even when ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})\simeq{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1})$, the approximation error from the CFG is given by
$\boldsymbol{\varepsilon}_{cfg}$ $:={\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})-{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t-1})=({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1}))+\omega\big{(}\delta{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-\delta{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t-1})\big{)}$
$\simeq\omega\big{(}\delta{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-\delta{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t-1})\big{)},$ (26)
where $\delta{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t}):={\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})$ denotes the directional component from unconditional to the conditional diffusion. Since the directional component by the text guidance is not negligible, this error becomes significant for high guidance scale $\omega$. Accordingly, the guidance scale must be heavily downweighted in order for inversions on real world images to be stable, thus limiting the strength of edits. To mitigate this issue, the authors in (mokady2023null; wallace2023edict) developed null text optimization and coupled transform techniques, respectively.
On the other hand, under the usual DDIM assumption ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})\simeq{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1})$, the approximation error of CFG++ mainly arises from (25), which is smaller than that of CFG since we have
$\|\boldsymbol{\varepsilon}_{cfg++}\|=\lambda\|\delta{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-\delta{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t-1})\|<\|\boldsymbol{\varepsilon}_{cfg}\|$ (27)
thanks to $\lambda<\omega$. Therefore, CFG++ significantly improves the DDIM inversion as shown Fig. 1 (middle) for representative results and Sec. 4.2 for further discussions.
Figure 7: T2I using SD v1.5, CFG vs CFG++ ($\omega=9.0,\lambda=0.8$). Unnatural depictions of human hands, and incorrect renderings of the text by CFG are corrected in CFG++
$\omega=2.0,\lambda=0.2$ $\omega=5.0,\lambda=0.4$ $\omega=7.5,\lambda=0.6$ $\omega=9.0,\lambda=0.8$ $\omega=12.5,\lambda=1.0$
Method FID $\downarrow$ CLIP $\uparrow$ FID $\downarrow$ CLIP $\downarrow$ FID $\downarrow$ CLIP $\uparrow$ FID $\downarrow$ CLIP $\uparrow$ FID $\downarrow$ CLIP $\uparrow$
CFG (ho2021classifierfree) 13.84 0.298 15.08 0.310 17.71 0.312 20.01 0.312 21.23 0.313
CFG++ (ours) 12.75 0.303 14.95 0.310 17.47 0.312 19.34 0.313 20.88 0.313
Table 1: Quantitative evaluation (FID, CLIP-similarity) of T2I with SD v1.5
4 Experimental Results
In this section, we design experiments to show the limitation of CFG and how CFG++ can effectively mitigate these downsides. We select Stable Diffusion (rombach2022high) (SD) as our baseline model for its widespread use. Experiments are conducted using SD v1.5 from diffusers (patil2022stable) unless specified otherwise. Notice that we use the conditioning scale $\lambda\in[0,1]$ for CFG++ whereas for standard CFG, the typical range is $\omega>1.0$. When searching for the matched guidance values of $\omega$ and $\lambda$ for a fair comparison, we fix $\lambda=0.2,0.4,0.6,0.8,1.0$ and find the $\omega$ values that produce the images that are of closest proximity in terms of LPIPS distance given the same seed. We found that the corresponding values were $\omega=2.0,5.0,7.5,9.0,12.5$, respectively, as shown in Appendix Fig. 14.
4.1 Text-to-Image Generation
Using the corresponding scales for $\omega$ and $\lambda$, we directly compare the performance of the T2I task using SD v1.5 and SDXL. In Tab. 1, we observe a constant improvement of the FID metric across all guidance scales, with approximately the same level of CLIP similarity or better. The improvements can also be clearly seen in Fig. 7 (SD v1.5), where the unnatural components of the generated images are corrected. Specifically, we see that unnatural depictions of human hands, and incorrect renderings of the text are corrected in CFG++, a long-standing research question in and of its own (podell2023sdxl; pelykh2024giving; chen2023textdiffuser).
Figure 8: Interpolating behavior of T2I along various guidance scales using SDXL with CFG++. The generated visuals align more significantly with the given textual description as the guidance scale increase to 1.
In Fig. 8, we visualize multiple images generated by CFG++ as we increase the guidance scale $\lambda$, illustrating a smooth transition from unconditional to highly conditional sampling. Specifically, at higher guidance scales, the target concept is more pronounced, with less emphasis on noisy, unrelated elements. This results in images that more closely reflect the specified text conditions. For example, in the second row, the images increasingly resemble a "painting by Van Gogh," demonstrating how the generated visuals align more significantly with the given textual description as the guidance scale increase to 1.
4.2 Diffusion image Inversion and Editing
Figure 9: Real image DDIM inversion results using SDv1.5. (a) Reconstructed samples after inversion by CFG and CFG++. (b) Quantitative comparison between CFG and CFG++.
Figure 10: Real image inversion results at various CFG and CFG++ scales using SDv1.5. The image is from COCO dataset. We use the text prompt “a white dog is sleeping on a street and a bicycle” during the inversion.
We further explore the effect of CFG++ on DDIM inversion (dhariwal2021diffusion), where the source image is reverted to a latent vector that can reproduce the original image through generative process. DDIM inversion is well-known to break down in the usual CFG setting as CFG magnifies the accumulated error of each inversion step (mokady2023null), violating local linearization assumptions (wallace2023edict). Remarkably, CFG++ may mitigate this issue. This improves inversion and enables better image editing compared to inversion with the original CFG. We evaluate our method following the experimental setups in park2024energy and kim2024dreamsampler. More details follow.
Diffusion Image Inversion. Using the matched set of scales for $\omega$ and $\lambda$, we demonstrate the effect of CFG++ on the diffusion image inversion task. Specifically, we reconstruct the images after inversion and evaluate it through PSNR and RMSE, following the methodology of wallace2023edict. In Fig. 9, we illustrate the reconstructed examples (9a) from a real image from COCO data set and computed metrics (9b) for 5k COCO-2014 (lin2014microsoft) validation set. Both qualitative and quantitative evaluation demonstrate a consistent improvement on reconstruction performance induced by CFG++. Notably, DDIM inversion with CFG++ leads to consistent reconstruction of the source image across all guidance scales, while DDIM inversion with CFG fails to reconstruct it, as shown in Fig. 10 for another real image.
Image Editing. Fig. 11 compares the image editing result using CFG and CFG++ followed by image inversion. In the image editing stage, a word in the source text on the left side is swapped with the target concept, and this modified text is used as condition for sampling. For both synthetic and real images, CFG++ successfully edits the target concept while preserving other concepts, such as background. In particular, the water splashing or brush in the background, which tend to disappear in the conventional CFG, are maintained through the inversion process by CFG++. This emphasizes CFG++’s superior ability to retain specific scene elements that are frequently lost in the standard CFG approach.
Figure 11: Comparison on image editing via DDIM inversion using SDXL: (top) synthetic image editing, (bottom) real image editing.
4.3 Text-Conditioned Inverse Problems
Inverse problem involves restoring the original data ${\boldsymbol{x}}$ from a noisy measurement ${\boldsymbol{y}}={\mathcal{A}}({\boldsymbol{x}})+{\boldsymbol{n}}$, where ${\mathcal{A}}$ represents an imaging operator introduces distortions (e.g, Gaussian blur) and ${\boldsymbol{n}}$ denotes the measurement noise. Diffusion inverse solvers (DIS) address this challenge by leveraging pre-trained diffusion models as implicit priors and performing posterior sampling ${\boldsymbol{x}}\sim p({\boldsymbol{x}}|{\boldsymbol{y}})$. Methods that leverage latent diffusion have gained recent interest (rout2024solving; song2024solving), but leveraging texts for solving these problems remains relatively underexplored (chung2023prompt; kim2023regularization). One of the main reasons for this is that naively using CFG on top of latent DIS leads to diverging samples (chung2023prompt). Several heuristic modifications such as null prompt optimization (kim2023regularization) with modified sampling schemes were needed to mitigate this drawback. This naturally leads to the question: Is it possible to leverage CFG guidance as a plug-and-play component of existing solvers? Here, we answer this question with a positive by showing that CFG++ enables the incorporation of text prompts into a standard solver. Specifically, we focus on comparing the performance of PSLD (rout2024solving) combined with CFG and CFG++ in solving linear inverse problems. This evaluation utilizes the FFHQ (karras2019style) 512x512 dataset and the text prompt "a high-quality photo of a face". Further details about the experimental settings can be found in the Appendix A.
As shown in Tab. 2, our method mostly outperforms both the vanilla PSLD and PSLD with CFG. The superiority is also evident from the Fig. 12, where CFG++ consistently delivers high-quality reconstructions across all tasks. PSLD with CFG often suffer from artifacts and blurriness. Conversely, CFG++ achieves better fidelity, clearly distinguishing between faces and faithfully reproducing fine details like eyelids and hair texture. For more results, please refer to the Appendix B.
Figure 12: Qualitative comparison on various inverse problems using PSLD (rout2024solving) under CFG and CFG++.
SR (x8) Deblur (motion) Deblur (gauss) Inpaint
Method FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$ FID $\downarrow$ LPIPS $\downarrow$ PSNR $\uparrow$
PSLD 46.24 0.413 24.41 97.51 0.500 21.83 41.65 0.388 26.88 10.27 9.36 30.15
PSLD + CFG 41.24 0.394 24.91 91.90 0.493 22.29 41.52 0.390 26.94 9.36 0.055 30.27
PSLD + CFG++ (ours) 36.58 0.385 24.87 65.67 0.482 21.93 39.85 0.400 26.90 9.78 0.052 30.31
Table 2: Quantitative comparison (FID, LPIPS, PSNR) of PSLD, PSLD with CFG, and PSLD with CFG++ on Latent Diffusion Inverse Solver.
5 Related Works and Discussion
Related works. kynkaanniemi2024applying observed that applying CFG guidance at the earlier stages of sampling always led to detrimental effects and drastically reduced the diversity of the samples. The guidance at the later stages of sampling had minimal effects. Drawing upon these observations, it was empirically shown that applying CFG only in the limited time interval near the middle led to the best performance. Similar observations were made in the SD community (howard2022; birch2023) in adjusting the guidance scale across $t$. These findings are orthogonal to ours in that these methods keep the sampling trajectory the same and try to empirically adjust the strength of the guidance, while we aim to design a different trajectory.
Implications. In several studies, attempts at designing a better trajectory of the generative path were made (liu2023flow; lipman2023flow). As diffusion models can only estimate the tangent vector of the trajectory at time $t$, it is now a well regarded consensus that minimizing the curvature of the trajectory leads to faster sampling (liu2023flow; lee2023minimizing) and better distillation results (liu2023instaflow). In this regard, the excessively high curvature and oscillating behavior of CFG with high guidance scales would hamper the performance of fast sampling and distillation, which was constantly reported in the literature as difficulties in extending distillation methods (salimans2022progressive; song2023consistency) to the latent space (meng2023distillation; luo2023latent; chen2024pixart). Straightening the diffusion path prior to distillation with Reflow (liu2023flow) exhibited much better results than aiming for direct distillation. In this regard, CFG++ can also be considered a better alternative to CFG for enhanced distillation performance.
6 Conclusion
This paper revises the most widespread guidance method, classifier-free guidance (CFG), highlighting that a small and reasonable guidance scale, e.g. $\lambda\in[0.0,1.0]$, might suffice successful guidance with a proper geometric correction. Observing that the original CFG suffers from off-manifold issues during sampling, we propose a simple but surprisingly effective fix: using ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})$ instead of ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$ in updating ${\boldsymbol{x}}_{t-1}$. This change translates the conditional guidance from extrapolating to interpolating between unconditional and conditional sampling trajectories, leading to more interpretable guidance in contrast to the conventional CFG which heavily relies on heuristic $\omega$ scaling. Given that CFG++ mitigates off-manifold issues, it may be beneficial for other downstream applications requiring accurate latent denoised estimate representations, e.g. DIS, as demonstrated in our experiments.
Appendix A Experimental details
A.1 Text-Conditioned Inverse Problems
Problem settings. We evaluate our approach across following degradation types: 1) Super-resolution with a scale factor of x8, 2) Motion deblurring from an image convolved with a 61 x 61 motion kernel, randomly sampled with an intensity value $0.3^{2}$, 3) Gaussian deblurring from an image convolved with a 61 x 61 kernel with an intensity value 0.5, and 4) Inpainting from 10-20% free-form masking, as implemented in (saharia2022palette). In inpainting evaluations, regions outside the mask are overlaid with the ground truth image.
Datasets, Models. For evaluation, we use the FFHQ (karras2019style) 512x512 dataset and follow (chung2023diffusion) by selecting the first 1,000 images for testing. For the pre-trained latent diffusion model including baseline methods, we choose SD v1.5 trained on the LAION dataset. As a baseline for latent DIS, we consider PSLD (rout2024solving). It enforces fidelity by projecting onto the subspace of ${\mathcal{A}}$ during the intermediate step between decoding and encoding, in conjunction with the DPS (chung2023diffusion) loss term. For gradient updates in both vanilla PSLD and PSLD with CFG, we use static step sizes of $\eta=1.0$ and $\gamma=0.1$ as recommended in (rout2024solving). For CFG scale $\omega$, we applied the corresponing scales that we found to be corresponded for CFG++ scale $\lambda$. Please refer to the Tab. 3 for the hyperparameters used for PSLD with CFG++.
SR(x8) Deblur(motion) Deblur(gauss) Inpaint
$\eta$ 1.3 0.4 1.0 1.0
$\gamma$ 0.1 0.025 0.12 0.1
$\lambda$ 0.1 0.2 0.2 0.6
$\omega$ 1.5 2.0 2.0 7.5
Table 3: Hyperparameters for PSLD (rout2024solving) with CFG++ and corresponding CFG scale $\omega$.
Appendix B Further experimental results
B.1 T2I
Since the range of guidance scales for CFG and CFG++ is different, we matched the guidance values of $\omega$ and $\lambda$ for comparison by computing the LPIPS distance using the same seed. Fig. 14 displays an example of generated images given the same text prompt and seed, while varying the guidance scale.
Fig. 13 shows the clean estimates at each diffusion sampling time $t$, computed using Tweedie’s formula, both with and without CFG. It is evident that the original CFG induces significant error, particularly at earlier times, leading to off-manifold samples. However, the CFG++ addresses this issue by adjusting the DDIM re-noising step and the guidance scale.
Figure 13: The discrete evolution of the posterior mean in CFG and CFG++. Denoised estimates in latent space are decoded into pixel space at each timestep using SDXL.
Figure 14: Example of generated images with various guidance scales by CFG and CFG++. The respective guidance scales are chosen to best match the generated images by CFG and CFG++.
B.2 Real Image Editing
In Fig. 15-17, we provide qualitative comparison on real image editing via DDIM inversion with CFG and CFG++. For the DDIM inversion stage, we use "a photography of [source concept]" as conditioning prompt. For the sampling stage, we swap [source concept] to [target concept] and use it as conditioning prompt for generation. For example, in Fig. 15, we set [source concept] to "dog" and [target concept] to "cat". For all experiments, we set the guidance scale as $\omega=9.0$ and $\lambda=0.8$ as described in the main paper. The comparison demonstrates that CFG++ successfully edits the given image which was not possible by CFG. This results also support our claim on reduced error during DDIM inversion by CFG++.
Figure 15: Comparison on real image editing via DDIM inversion under CFG and CFG++ using SDXL. Cat $\rightarrow$ Dog.
Figure 16: Comparison on real image editing via DDIM inversion under CFG and CFG++ using SDXL. Horse $\rightarrow$ Zebra.
Figure 17: Comparison on real image editing via DDIM inversion under CFG and CFG++ using SDXL. Cat $\rightarrow$ Cat with glasses.
B.3 Text-Conditioned Inverse Problems
In Fig. 18-21, we display additional qualitative comparison for text-conditioned inverse problem solver with CFG and CFG++. Experiments are conducted with FFHQ (512x512) validation set and CFG++ consistently leads to better reconstruction of true solution for various tasks.
Figure 18: Results of PSLD using CFG and CFG++ on the FFHQ dataset at Super-Resolution (x8).
Figure 19: Results of PSLD using CFG and CFG++ on the FFHQ dataset at Motion Deblurring.
Figure 20: Results of PSLD using CFG and CFG++ on the FFHQ dataset at Gaussian Deblurring.
Figure 21: Results of PSLD using CFG and CFG++ on the FFHQ dataset at Inpainting.

### v2

@@
 Contrary to the widespread belief that these are inherent limitations of diffusion models, this paper reveals that the problems actually stem from the off-manifold phenomenon associated with CFG, rather than the diffusion models themselves.
-More specifically, inspired by the recent advancements of diffusion model-based inverse problem solvers (DIS), we reformulate text-guidance as an inverse problem with a text-conditioned score matching loss, and develop CFG++, a novel approach that tackles the off-manifold challenges inherent in traditional CFG.
+More specifically, inspired by the recent advancements of diffusion model-based inverse problem solvers (DIS), we reformulate text-guidance as an inverse problem with a text-conditioned score matching loss and develop CFG++, a novel approach that tackles the off-manifold challenges inherent in traditional CFG.
 CFG++ features a surprisingly simple fix to CFG, yet it offers significant improvements, including better sample quality for text-to-image generation, invertibility, smaller guidance scales, reduced mode collapse, etc.
 Furthermore, CFG++ enables seamless interpolation between unconditional and conditional sampling at lower guidance scales, consistently outperforming traditional CFG at all scales.
+Moreover, CFG++ can be easily integrated into the high-order diffusion solvers and naturally extends to distilled diffusion models.
 Experimental results confirm that our method significantly enhances performance in text-to-image generation, DDIM inversion, editing, and solving inverse problems, suggesting a wide-ranging impact and potential applications in various fields that utilize text guidance.
 Project Page: https://cfgpp-diffusion.github.io/.
-Figure 1: (Top) Comparison of T2I results by SDXL for the prompt "kayak in the water, optical color, aerial view, rainbow".
-The CFG-guided image has a broken oar, an artifact significantly reduced in the CFG++ version.
+Figure 1: (Top) Comparison of T2I results by SDXL-Lightning for the prompt "kayak in the water, optical color, aerial view, rainbow".
+The CFG-guided image has significant artifacts, which are reduced in the CFG++ version.
 (Middle) DDIM Inversion results under CFG show noticeable artifacts at various CFG scales, which are significantly reduced by CFG++.
 (Bottom) The evolution of denoised estimates differs between CFG and CFG++.
-CFG exhibits sudden shifts and intense color saturation early in reverse diffusion, while CFG++ transitions smoothly from low to high-resolution details.
+CFG exhibits sudden shifts and intense color saturation early in reverse diffusion, while CFG++ transitions smoothly from low to high-resolution.
 1 Introduction
@@
 To emphasize the guidance, one uses the guidance scale $\omega>1$, where the direction can be defined by the direction from the unconditional score to the conditional score (Ho & Salimans, 2021).
-In modern text-to-image (T2I) diffusion models, the guidance scale $\omega$ is typically set within the range of $[5.0,30]$, referred to as the moderately high range of CFG guidance  (Chen et al., 2024b; Podell et al., 2023).
+In modern text-to-image (T2I) diffusion models, the guidance scale $\omega$ is typically set within the range of $[5.0,30]$, referred to as the moderately high range of CFG guidance  (Chen et al., 2024; Podell et al., 2023).
 The insufficiency in guidance also holds for classifier guidance (Dhariwal & Nichol, 2021; Song et al., 2021b) so that a scale of 10 was used.
@@
 In this work, we aim to give an answer to this conundrum by revisiting the geometric view of diffusion models.
-In particular, inspired by the recent advances in diffusion-based inverse problem solvers (DIS) (Kadkhodaie & Simoncelli, 2021; Chung et al., 2023a; Song et al., 2023a; Kim et al., 2024; Chung et al., 2024), we reformulate the text-guidance as an inverse problem with a text-conditioned score matching loss and derive reverse diffusion sampling strategy by utilizing the decomposed diffusion sampling (DDS) (Chung et al., 2024).
+In particular, inspired by the recent advances in diffusion-based inverse problem solvers (DIS) (Kadkhodaie & Simoncelli, 2021; Chung et al., 2023a; Song et al., 2023; Kim et al., 2024; Chung et al., 2024), we reformulate the text guidance as an inverse problem with a text-conditioned score-matching loss and derive a reverse diffusion sampling strategy by utilizing decomposed diffusion sampling (DDS) (Chung et al., 2024).
 This results in a surprisingly simple fix of CFG to the sampling process without any computational overhead.
@@
 Sampling from the data distribution can be performed by solving either the reverse stochastic differential equation (SDE) or the equivalent probability-flow ordinary differential equation (PF-ODE) (Song et al., 2021b).
-For example, under the choice $p({\boldsymbol{x}}_{t}|{\boldsymbol{x}}_{0})=\mathcal{N}({\boldsymbol{x}}_{0},t^{2}{\bm{I}})$, the generative PF-ODE reads
-$d{\boldsymbol{x}}_{t}=-t\nabla_{{\boldsymbol{x}}_{t}}\log p({\boldsymbol{x}}_{t})\,dt=\frac{{\boldsymbol{x}}_{t}-\mathbb{E}[{\boldsymbol{x}}_{0}|{\boldsymbol{x}}_{t}]}{t}\,dt,\,\,{\boldsymbol{x}}_{T}\sim p_{T}({\boldsymbol{x}}_{T}),$ (1)
-where Tweedie’s formula (Efron, 2011) ${\mathbb{E}}[{\boldsymbol{x}}_{0}|{\boldsymbol{x}}_{t}]={\boldsymbol{x}}_{t}+t^{2}\nabla_{{\boldsymbol{x}}_{t}}\log p({\boldsymbol{x}}_{t})$ is applied to achieve the second equality.
-Thus, the score function $\nabla_{{\boldsymbol{x}}_{t}}\log p({\boldsymbol{x}}_{t})$ plays the key role in diffusion models.
-One can approximate an equivalent parameterization of the score function via denoising score matching (Hyvärinen & Dayan, 2005) or epsilon-matching (Ho et al., 2020), etc.
-When aiming for a text-conditional diffusion model that can condition on arbitrary ${\boldsymbol{c}}$, one can extend epsilon matching to a conditional one as:
-$\min_{\theta}{\mathbb{E}}_{{\boldsymbol{x}}_{t}\sim p({\boldsymbol{x}}_{t}|{\boldsymbol{x}}_{0}),\{{\boldsymbol{x}}_{0},{\boldsymbol{c}}\}\sim p({\boldsymbol{x}}_{0},{\boldsymbol{c}}),{\boldsymbol{\epsilon}}\sim{\mathcal{N}}(0,{\bm{I}})}\left[\|{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},{\boldsymbol{c}})-{\boldsymbol{\epsilon}}\|_{2}^{2}\right],$ (2)
-where the condition is dropped with a certain probability (Ho & Salimans, 2021) so that null conditioning with ${\boldsymbol{c}}=\varnothing$ is possible.
+For example, under the choice $p({\boldsymbol{x}}_{t}|{\boldsymbol{x}}_{0})=\mathcal{N}({\boldsymbol{x}}_{0},\sigma_{t}^{2}{\bm{I}})$, the generative PF-ODE reads
+$d{\boldsymbol{x}}_{t}=-\sigma_{t}\nabla_{{\boldsymbol{x}}_{t}}\log p({\boldsymbol{x}}_{t})\,dt=\frac{{\boldsymbol{x}}_{t}-\mathbb{E}[{\boldsymbol{x}}_{0}|{\boldsymbol{x}}_{t}]}{\sigma_{t}}\,dt,\,\,\quad{\boldsymbol{x}}_{T}\sim p_{T}({\boldsymbol{x}}_{T}),$ (1)
+where Tweedie’s formula (Efron, 2011) ${\mathbb{E}}[{\boldsymbol{x}}_{0}|{\boldsymbol{x}}_{t}]={\boldsymbol{x}}_{t}+\sigma_{t}^{2}\nabla_{{\boldsymbol{x}}_{t}}\log p({\boldsymbol{x}}_{t})$ is applied to achieve the second equality.
+When aiming for a text-conditional diffusion model that can condition on arbitrary ${\boldsymbol{c}}$, one can extend epsilon matching to a conditional one where the condition is dropped with a certain probability (Ho & Salimans, 2021) so that null conditioning with ${\boldsymbol{c}}=\varnothing$ is possible.
 The neural network architecture of ${\boldsymbol{\epsilon}}_{\theta}$ is designed so that the condition ${\boldsymbol{c}}$ can modulate the output through cross attention (Rombach et al., 2022).
 For simplicity in notation throughout the paper, we define ${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}:={\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},{\boldsymbol{c}})$ and ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}:={\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},\varnothing)$ by dropping $\theta$ and ${\boldsymbol{x}}_{t}$.
-Extending the result of the Tweedie’s formula to the unconditional case under the variance preserving (VP) framework of DDPMs (Ho et al., 2020), we have
+Extending the result of Tweedie’s formula to the unconditional case under the variance preserving (VP) framework of DDPMs (Ho et al., 2020), we have
 ${\mathbb{E}}[{\boldsymbol{x}}_{0}|{\boldsymbol{x}}_{t},\varnothing]=\hat{\boldsymbol{x}}_{\varnothing}({\boldsymbol{x}}_{t}):=({\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t}))/\sqrt{\bar{\alpha}_{t}}.$ (2)
@@
 ${\boldsymbol{x}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}},$ (4)
-where ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}:=\hat{\boldsymbol{x}}_{\varnothing}({\boldsymbol{x}}_{t})={\mathbb{E}}[{\boldsymbol{x}}_{0}|{\boldsymbol{x}}_{t},\varnothing]$ is the denosed signal by Tweedie’s formula, and (5) corresponds to the renoising step.
+where ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}:=\hat{\boldsymbol{x}}_{\varnothing}({\boldsymbol{x}}_{t})={\mathbb{E}}[{\boldsymbol{x}}_{0}|{\boldsymbol{x}}_{t},\varnothing]$ is the denoised signal by Tweedie’s formula, and (4) corresponds to the renoising step.
 This is repeated for $t=T,T-1,\dots,1$.
@@
 Diffusion model-based inverse problem solvers.
-Diffusion model-based inverse problem solvers (DIS) aim to perform posterior sampling from an unconditional diffusion model (kadkhodaie2021stochastic; chung2023diffusion; song2023pseudoinverseguided; kim2024dreamsampler).
-Specifically, for a given loss function $\ell({\boldsymbol{x}})$ which often stems from the likelihood for measurement consistency [^1] For example, for a given measurement model ${\boldsymbol{y}}=A{\boldsymbol{x}}+{\boldsymbol{w}}$ with ${\boldsymbol{w}}\sim\mathcal{N}({\boldsymbol{x}}_{0},\sigma^{2}{\bm{I}})$, the loss function is the data consistency (DC) loss given by $\ell({\boldsymbol{x}})=\|{\boldsymbol{y}}-A{\boldsymbol{x}}\|^{2}/2\sigma^{2}$.
-, the goal of DIS is to address the following optimization problem
-$\min_{{\boldsymbol{x}}\in{\mathcal{M}}}\ell({\boldsymbol{x}})$ (8)
-where ${\mathcal{M}}$ represents the clean data manifold sampled from unconditional distribution $p_{0}({\boldsymbol{x}})$.
-Consequently, it is essential to navigate in a way that minimizes cost while also identifying the correct clean manifold.
-Recently, chung2023diffusion proposed a general technique called diffusion posterior sampling (DPS), where the updated estimate from the noisy sample ${\boldsymbol{x}}_{t}\in{\mathcal{M}}_{t}$ is constrained to stay on the same noisy manifold ${\mathcal{M}}_{t}$.
-This is achieved by computing the manifold constrained gradient (MCG) (chung2022improving) on a noisy sample ${\boldsymbol{x}}_{t}\in{\mathcal{M}}_{t}$ as $\nabla^{mcg}_{{\boldsymbol{x}}_{t}}\ell({\boldsymbol{x}}_{t}):=\nabla_{{\boldsymbol{x}}_{t}}\ell(\hat{\boldsymbol{x}}_{t})$, where $\hat{\boldsymbol{x}}_{t}$ is the denoised sample in (3) through Tweedie’s formula.
-The resulting algorithm can be stated as follows:
+Diffusion model-based inverse problem solvers (DIS) aims to perform posterior sampling from an unconditional diffusion model (kadkhodaie2021stochastic; chung2023diffusion; song2023pseudoinverseguided; kim2024dreamsampler).
+Specifically, for a given loss function $\ell({\boldsymbol{x}})$ which often stems from the likelihood, the goal of DIS is to address the optimization problem $\min_{{\boldsymbol{x}}\in{\mathcal{M}}}\ell({\boldsymbol{x}})$, where ${\mathcal{M}}$ represents the clean data manifold sampled from the unconditional distribution $p_{0}({\boldsymbol{x}})$.
+It is essential to navigate in a way that minimizes cost while also identifying the correct clean manifold.
+chung2023diffusion proposed diffusion posterior sampling (DPS), where the updated estimate from the noisy sample ${\boldsymbol{x}}_{t}\in{\mathcal{M}}_{t}$ is constrained to stay on the same noisy manifold ${\mathcal{M}}_{t}$.
+This is achieved by computing the manifold constrained gradient (MCG) (chung2022improving) on a noisy sample ${\boldsymbol{x}}_{t}\in{\mathcal{M}}_{t}$ as $\nabla^{mcg}_{{\boldsymbol{x}}_{t}}\ell({\boldsymbol{x}}_{t}):=\nabla_{{\boldsymbol{x}}_{t}}\ell(\hat{\boldsymbol{x}}_{t})$, where $\hat{\boldsymbol{x}}_{t}$ is the denoised sample in (2) through Tweedie’s formula (efron2011tweedie).
+The resulting algorithm with DDIM (song2020score) can be stated as follows:
 ${\boldsymbol{x}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}\left({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}-\gamma_{t}\nabla_{{\boldsymbol{x}}_{t}}\ell({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})\right)+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}},$ (7)
@@
 Under the linear manifold assumption (chung2022improving; chung2023diffusion), this allows precise transition to ${\mathcal{M}}_{t-1}$.
-Therefore, by performing (9) from $t=T$ to $t=0$, we can solve the optimization problem (8) with ${\boldsymbol{x}}_{0}\in{\mathcal{M}}$.
-Unfortunately, the computation of MCG requires computationally expensive backpropagation and is often unstable  (poole2022dreamfusion; du2023reduce).
-In another work, chung2024decomposed shows that (9) can be equivalently represented as
-${\boldsymbol{x}}_{t-1}$ $=\sqrt{\bar{\alpha}_{t-1}}P_{\mathcal{M}}\left({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}-\gamma_{t}\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})\right)+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}},$ (10)
-where $P_{\mathcal{M}}$ denotes the projection on the clean manifold ${\mathcal{M}}$.
-This implies that the data consistency update step becomes a standard gradient step, followed by the projection on the clean manifold.
-Importantly, under the commonly used linear manifold assumption in DPS, the one step update by ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}-\gamma_{t}\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})$ or multi-step update employing the conjugate gradient (CG) method are guaranteed to remain within a linear subspace, thus obviating the need for explicit computation of $P_{\mathcal{M}}$, i.e.
-${\boldsymbol{x}}_{t-1}$ $\simeq\sqrt{\bar{\alpha}_{t-1}}\left({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}-\gamma_{t}\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})\right)+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}.$ (11)
-This method, often referred to as the decomposed diffusion sampling (DDS), bypasses the computation of the score Jacobian, similar to poole2022dreamfusion, making it stable and suitable for large scale medical imaging inverse problems (chung2024decomposed).
+To mitigate the computational complexity and the instability of neural network backprop, chung2024decomposed shows that (7) can be equivalently represented as
+${\boldsymbol{x}}_{t-1}$ $\simeq\sqrt{\bar{\alpha}_{t-1}}\left({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}-\gamma_{t}\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})\right)+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}$ (8)
+under further assumptions on ${\mathcal{M}}$.
+This method, often referred to as the decomposed diffusion sampling (DDS), bypasses the computation of the score Jacobian, similar to poole2022dreamfusion, making it stable and suitable for large-scale medical imaging inverse problems (chung2024decomposed).
 In the following, we leverage the insight from DDS to propose an algorithm to improve upon the CFG algorithm.
@@
 3.1 Derivation of CFG++
-Instead of uncritically adopting the sharpened posterior distribution $p^{\omega}({\boldsymbol{x}}|{\boldsymbol{c}})\propto p({\boldsymbol{x}})p({\boldsymbol{c}}|{\boldsymbol{x}})^{\omega}$ as introduced by ho2021classifierfree and further explored in the Bayesian CFG formulation, we adopt a fundamentally different strategy by reformulating text-guidance as an optimization problem.
-Specifically, our focus is on identifying a loss function $\ell({\boldsymbol{x}})$ in (8) such that, when minimized under the condition set by the text, enables the reverse diffusion process to progressively generate samples that increasingly satisfy the text condition.
-One of the most significant contributions of this paper is to reveal that the text-conditioned score matching loss or score distillation sampling (SDS) loss is also ideally suited for our purpose.
+Instead of uncritically adopting the sharpened posterior distribution $p^{\omega}({\boldsymbol{x}}|{\boldsymbol{c}})\propto p({\boldsymbol{x}})p({\boldsymbol{c}}|{\boldsymbol{x}})^{\omega}$ as introduced by ho2021classifierfree, we adopt a fundamentally different strategy by reformulating text-guidance as an optimization problem.
+Specifically, our focus is on identifying a loss function $\ell({\boldsymbol{x}})$ in (8) such that, when minimized under the condition set by the text, enables the reverse diffusion process to generate samples that increasingly satisfy the text condition progressively.
+One of the most significant contributions of this paper is to reveal that the text-conditioned score matching loss or alternatively, score distillation sampling (SDS) loss (poole2022dreamfusion) is ideally suited for our purpose.
 Specifically, we are interested in solving the following inverse problem through diffusion models:
 $\min_{{\boldsymbol{x}}\in{\mathcal{M}}}\ell_{sds}(x),\quad\ell_{sds}({\boldsymbol{x}}):=\|{\boldsymbol{\epsilon}}_{\theta}(\sqrt{\bar{\alpha}_{t}}{\boldsymbol{x}}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}},{\boldsymbol{c}})-{\boldsymbol{\epsilon}}\|_{2}^{2}$ (9)
-This implies that our goal is to identify solutions on the clean manifold ${\mathcal{M}}$ that optimally align with the text condition ${\boldsymbol{c}}$.
-Figure 4: Enhanced T2I results by SDXL ($\omega=9.0,\lambda=0.8$) with CFG++.
-Under CFG, the lion cub is not visible (top-left), the dog appears with two tails (top-right), the goggles have an unusual shape (bottom-left), and the tree trunk is folded (bottom-right).
-These artifacts are absent in those produced by CFG++.
-To avoid the Jacobian computation, in this paper we attempt to solve (12) through decomposed diffusion sampling (DDS). The resulting sampling process from reverse diffusion is then given by
-${\boldsymbol{x}}_{t-1}$ $=$ $\sqrt{\bar{\alpha}_{t-1}}\left({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}-\gamma_{t}\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell_{sds}({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})\right)+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}$ (13)
-which is analogous to (11).
-Consequently, the remaining technical issue is to compute $\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell_{sds}({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})$.
-To answer this, by using the forward diffusion to generate ${\boldsymbol{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\boldsymbol{x}}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}$ from the clean image ${\boldsymbol{x}}\in{\mathcal{M}}$, we convert the objective function in (12) into the following equivalent loss:
-$\ell_{sds}({\boldsymbol{x}})$ $=\|{\boldsymbol{\epsilon}}-{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},{\boldsymbol{c}})\|^{2}=\left\|\frac{{\boldsymbol{x}}_{t}-\sqrt{\bar{\alpha}_{t}}{\boldsymbol{x}}}{\sqrt{1-\bar{\alpha}_{t}}}-\frac{{\boldsymbol{x}}_{t}-\sqrt{\bar{\alpha}_{t}}{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}}{\sqrt{1-\bar{\alpha}_{t}}}\right\|^{2}$ (14)
-$=\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}\|{\boldsymbol{x}}-{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}\|^{2}$ (15)
-where we use the conditional Tweedie formula derived by replacing ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}$ in (4) by the conditioned score ${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}$.
-Similar observation has been made in DreamSampler (kim2024dreamsampler) and HiFA (zhu2023hifa).
-Consequently, $\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell_{sds}({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})$ can be readily computable and the resulting DDS sampling process can be simplified as
+This implies that our goal is to identify solutions on the clean manifold ${\mathcal{M}}$ that optimally aligns with the text condition ${\boldsymbol{c}}$.
+To avoid the Jacobian computation, in this paper, we attempt to solve (9) through DDS in (8). The resulting sampling process from reverse diffusion is then given by
+${\boldsymbol{x}}_{t-1}$ $=$ $\sqrt{\bar{\alpha}_{t-1}}\left({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}-\gamma_{t}\nabla_{{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}}\ell_{sds}({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})\right)+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}.$ (10)
+By using ${\boldsymbol{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\boldsymbol{x}}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}$ from the clean image ${\boldsymbol{x}}\in{\mathcal{M}}$, we can equivalently write the loss as $\ell_{sds}({\boldsymbol{x}})=\frac{\bar{\alpha}_{t}}{1-\bar{\alpha}_{t}}\|{\boldsymbol{x}}-{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}\|^{2}$, which leads to
 ${\boldsymbol{x}}_{t-1}=\sqrt{\bar{\alpha}_{t-1}}\left({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}+\lambda({\hat{\boldsymbol{x}}_{\boldsymbol{c}}}-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})\right)+\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}$ (11)
 where $\lambda:=\frac{2{\bar{\alpha}_{t}}}{{1-\bar{\alpha}_{t}}}\gamma_{t}$.
-Although concise, (16) can be further simplified to a similar form of the traditional CFG as described in Algorithm 1.
-Specifically, note that
-${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}+\lambda({\hat{\boldsymbol{x}}_{\boldsymbol{c}}}-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})=({\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t}))/\sqrt{\bar{\alpha}_{t}}$ (17)
-where
-${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t}):={\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})+{\lambda}[{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})]$ (18)
-Therefore, (16) can be equivalently represented by:
+Using the CFG notation ${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t}):={\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})+{\lambda}[{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})]$ and ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}+\lambda({\hat{\boldsymbol{x}}_{\boldsymbol{c}}}-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}})=({\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t}))/\sqrt{\bar{\alpha}_{t}}$, (11) can be equivalently represented as
 ${\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})$ $=({\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t}))/\sqrt{\bar{\alpha}_{t}}$ (12)
@@
 which is summarized in Algorithm 2.
-By examining Algorithm 1 and Algorithm 2, we observe a notable similarity between CFG and CFG++.
-However, a crucial difference exists.
-Specifically, in CFG++, the renoising process after applying Tweedie’s formula should utilize the unconditional noise ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})$ instead of ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$.
-This surprisingly simple fix to the original CFG algorithm leads to smoother trajectory of generation, (Fig.
+By examining Algorithm 1 and Algorithm 2, we observe that CFG and CFG++ are mostly the same, with a crucial difference in the renoising process.
+This surprisingly simple fix of utilizing the unconditional noise ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})$ instead of ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$ leads to a smoother trajectory of generation, (Fig.
 1 bottom) and generation with superior quality (Fig.
 1 top).
-Moreover, CFG++ can also be easily extended to significantly improve DDIM inversion, which is known to be prone to failure by standard CFG (mokady2023null).
-See Fig.
-1 (middle) for representative results and Sec.
-4.2 for further discussions.
-In the following section, we present a theoretical analysis of CFG++ to investigate the origin of this improvement.
+Although we focus our construction of the solver on DDIM for simplicity, note that DDIM is just one way of reverse sampling.
+There are other widely-used solvers such as Karras Euler and its variants (karras2022elucidating), DPM-solver (lu2022dpm; lu2022dpmpp), their ancestral variants [^1] https://github.com/crowsonkb/k-diffusion, etc.
+For most widely used solvers up to the second order, a single-step update of solving the unconditional PF-ODE can be represented as
+${\boldsymbol{x}}_{i}={{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{i-1})}+a_{i}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{i-1})+b_{i}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{i-2})+c_{i}{\boldsymbol{x}}_{i-1}+d_{i}{\boldsymbol{\epsilon}},\,{\boldsymbol{\epsilon}}\sim{\mathcal{N}}(0,{\bm{I}}),$ (14)
+with $d_{i}\neq 0$ when one uses an ancestral sampler.
+Note that the first right-hand side term in (14) corresponds to the denosing, whereas the rest terms describes the higher-order corrected version of the renoising process.
+As the goal of CFG++ is to optimize the denoising process under the text-guidance while keeping the renoising components equivalent to unconditional sampling, applying CFG++ to the general iteration in (14) simply leads to
+${\boldsymbol{x}}_{i}={\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{i-1})+a_{i}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{i-1})+b_{i}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{i-2})+c_{i}{\boldsymbol{x}}_{i-1}+d_{i}{\boldsymbol{\epsilon}},\,{\boldsymbol{\epsilon}}\sim{\mathcal{N}}(0,{\bm{I}}).$ (15)
+Moreover, CFG++ naturally extends to distilled diffusion models, such as SDXL-turbo (sauer2023adversarial) and SDXL-lightning (lin2024sdxl), where akin to (15), we use the conditional denoised estimate, but for the rest of the noise components in the Euler solver, we use the unconditional estimate.
+For details in how to apply CFG++ to various solvers, see Appendix A.
 3.2 Geometry of CFG++
@@
 1 (bottom), we illustrate the evolution of the posterior mean through Tweedie’s formula during the reverse diffusion process.
-Notably, in the early phases of reverse diffusion sampling under CFG, there is a sudden shift in the image and intense color saturation, followed by only minor changes until later stages.
-This abrupt transition in the posterior mean is somewhat unexpected, as the reverse diffusion process is designed to progressively refine resolution, moving from a noisier to a clearer manifold.
-Similar phenomena are depicted in Fig.
-13 in Appendix.
-We refer to this as the off-manifold phenomenon, which will be clarified shortly.
-Conversely, under CFG++, the posterior mean evolves gradually, transitioning smoothly from a low-resolution image to high-resolution details.
-This indicates that CFG++ is less prone to the off-manifold phenomenon observed with CFG.
-In the following, we provide a theoretical analysis of the origin of the off-manifold phenomenon and explain why CFG++ can mitigate this issue.
-By examining (17) and (19) as well as Algorithm 1, we can easily see that the Tweedie’s formula for the CFG++ and CFG can be equivalently represented by
-${\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})=(1-\lambda){\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\lambda{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t}),\quad{\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})=(1-\omega){\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\omega{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})$ (21)
-This implies that $\lambda,\omega\in[0,1]$ faciliates an interpolation between the unconditional and the conditional posterior means.
-However, given the typical CFG scale $\omega>1.0$, we observe that CFG extrapolates beyond the unconditional and the conditional posterior means established by Tweedie’s formula.
-Consequently, under the usual assumption that the clean manifold to be piecewise linear around the posterior mean derived via Tweedie’s formula, the conditional posterior mean estimates from traditional CFG, obtained with a guidance scale outside the range of $[0,1]$, can extend beyond the piecewise linear manifold.
-This may lead to the estimates potentially ‘falling off’ the data manifold, as depicted by an orange arrow pointing downwards in Fig.
+Notably, in the early phases of reverse diffusion sampling under CFG, there is a sudden shift in the image and intense color saturation.
+Conversely, CFG++ is free from the undesirable off-manifold phenomenon.
+In the following, we investigate why this is the case.
+Note that the denoised estimate of CFG and CFG++ at time $t$ can be equivalently represented as
+${\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})=(1-\lambda){\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\lambda{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t}),\quad{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})=(1-\omega){\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\omega{\hat{\boldsymbol{x}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t}).$ (16)
+$\lambda,\omega\in[0,1]$ facilitates an interpolation.
+However, with $\omega>1.0$, CFG extrapolates beyond the unconditional and conditional estimates.
+Consequently, under the assumption that the clean manifold be piecewise linear (chung2024decomposed), the conditional posterior mean estimates from CFG obtained with a guidance scale outside the range of $[0,1]$, can easily extend beyond the piecewise linear manifold.
+This may lead to the estimates potentially “falling off” the data manifold, as depicted by an orange arrow pointing downwards in Fig.
 3(a).
-We conjecture that this is one of the primary contributors to the off-manifold phenomenon.
-Thus, we select $\lambda\in[0,1]$ as the guidance scale for CFG++ to ensure it remains an interpolation between the unconditional and conditional Tweedie, thus preventing it from ‘falling off’ the clean data manifold.
-An additional source of the off-manifold phenomenon in CFG occurs during the transition from the clean manifold ${\mathcal{M}}$ to the subsequent noisy manifold ${\mathcal{M}}_{t-1}$.
-Specifically, even from the same point in the clean manifold, the use of ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$ for this transition introduces a nonzero offset $\Delta^{\omega}:=\omega\sqrt{1-\bar{\alpha}_{t-1}}({\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t}))$ from ${\mathcal{M}}_{t-1}$, as illustrated in Fig.
+Thus, we select $\lambda\in[0,1]$ as the guidance scale for CFG++ to ensure it remains an interpolation between the unconditional and conditional estimate, thus preventing it from ‘falling off’ the clean data manifold.
+An additional source of the off-manifold phenomenon in CFG occurs during the transition from the clean manifold ${\mathcal{M}}$ to the subsequent noisy manifold ${\mathcal{M}}_{t-1}$, due to a similar off-manifold phenomenon from the large guidance scale (i.e.
+extrapolation), illustrated in Fig.
 3 (b).
-This leads to the invalid lifts from ${\mathcal{M}}$ to ${\mathcal{M}}_{t-1}$.
-On the other hand, we note that the renoising process of CFG++ follows the standard DDIM, resulting in the accurate transition to ${\mathcal{M}}_{t-1}$.
-Figure 5: The normalized text-conditioned score matching loss throughout the reverse diffusion sampling process for both CFG and CFG++.
+Figure 4: Text-conditioned score matching loss throughout the reverse diffusion sampling for both CFG and CFG++ in SDXL. Avg. loss computed with 55 prompts from (chen2024pixartalpha).
 Text-Image alignment.
 In Fig.
-1 (top), we display images generated by SDXL using the prompt “Kayak in the water, optical color, areal view, rainbow".
-The generated image on the left is guided by the traditional CFG, while the one on the right is produced with our CFG++.
-Although both images look impressive and appear to align well with the given text prompt, a closer examination reveals a significant difference.
-For instance, one of the oars in the CFG-guided image is broken in the middle, which shows a misalignment with the text prompt.
-This type of artifact is substantially reduced in the CFG++ version, where the oars are depicted intact.
-The examples provided by SDXL in Fig.
-4 further demonstrate the enhanced text-to-image alignment achieved with CFG++.
-For instance, the lion cub is not visible in Fig.
-4(top-left), the dog appears with two tails in Fig.
-4(top-right), the goggles have an unusual shape in Fig.
-4(bottom-left), and the tree trunk is folded in Fig.
-4(bottom-right).
-These artifacts, observed in the images generated by CFG, are absent in those produced by CFG++.
-In fact, this enhanced alignment capability is a natural consequence of CFG++, which directly minimizes the text-conditioned score-matching loss as shown in (12).
+1 (top), and Fig.
+9, we observe enhanced text-to-image alignment achieved with CFG++.
+This enhanced alignment capability is a natural consequence of CFG++, which directly minimizes the text-conditioned score-matching loss as shown in (9).
 In contrast, CFG indirectly seeks text alignment through the sharpened posterior distribution $p^{\omega}({\boldsymbol{x}}|{\boldsymbol{c}})\propto p({\boldsymbol{x}})p({\boldsymbol{c}}|{\boldsymbol{x}})^{\omega}$.
-Therefore, CFG++ inherently outperform CFG in terms of text alignment due to its fundamental design principle.
+Therefore, CFG++ inherently outperforms CFG in terms of text alignment due to its fundamental design principle.
 For example, Fig.
@@
 As discussed in (song2020denoising), the denoising process for unconditional DDIM is approximately invertible, meaning that ${\boldsymbol{x}}_{t}$ can generally be recovered from ${\boldsymbol{x}}_{t-1}$ Specifically, from (3) and (4), we have the following approximate inversion formula for unconditional DDIM:
-${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})$ $=({{\boldsymbol{x}}_{t-1}-\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})})/{\sqrt{\bar{\alpha}_{t-1}}}$
-$\simeq({{\boldsymbol{x}}_{t-1}-\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1})})/{\sqrt{\bar{\alpha}_{t-1}}}$ (22)
-${\boldsymbol{x}}_{t}$ $=\sqrt{\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})$
-$\simeq\sqrt{\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1})$ (23)
+${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})$ $=({{\boldsymbol{x}}_{t-1}-\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})})/{\sqrt{\bar{\alpha}_{t-1}}}\simeq({{\boldsymbol{x}}_{t-1}-\sqrt{1-\bar{\alpha}_{t-1}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1})})/{\sqrt{\bar{\alpha}_{t-1}}}$ (17)
+${\boldsymbol{x}}_{t}$ $=\sqrt{\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})\simeq\sqrt{\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1})$ (18)
 where the approximation arises from ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})\simeq{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1})$.
@@
 This formulation underpins the CFG++ guided DDIM inversion algorithm, presented in Algorithm 4.
-Algorithm 3 DDIM Inversion with CFG
-1:${\boldsymbol{x}}_{0},0\leq{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\omega}\in\mathbb{R}$
-2:for $i=0$ to $T-1$ do
-3:     ${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})={\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})+{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\omega}[{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})]$
-4:     ${\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})\leftarrow({\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t}))/\sqrt{\bar{\alpha}_{t}}$
-5:     ${\boldsymbol{x}}_{t+1}=\sqrt{\bar{\alpha}_{t+1}}{\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t+1}}{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$
-6:end for
-7:return ${\boldsymbol{x}}_{T}$
-Algorithm 4 DDIM Inversion with CFG++
-${\boldsymbol{x}}_{0},{\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\lambda}\in[0,1]$
-2:for $i=0$ to $T-1$ do
-3:     ${\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})={\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})+{\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\lambda}[{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})]$
-4:     ${\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})\leftarrow({\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t}))/\sqrt{\bar{\alpha}_{t}}$
-5:     ${\boldsymbol{x}}_{t+1}=\sqrt{\bar{\alpha}_{t+1}}{\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})+\sqrt{1-\bar{\alpha}_{t+1}}{\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})$
-6:end for
-7:return ${\boldsymbol{x}}_{T}$
-Figure 6: Comparison between DDIM inversion.
-CFG++ proposes a simple yet effective fix: using ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})$ instead of ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$ in Tweedie’s denoising step.
 In practice, for small time step size the error ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})\simeq{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t-1})$ is relatively small, so the unconditional DDIM inversion by (17) and (18) lead to relatively insignificant errors.
@@
 thanks to $\lambda<\omega$. Therefore, CFG++ significantly improves the DDIM inversion as shown Fig. 1 (middle) for representative results and Sec. 4.2 for further discussions.
-Figure 7: T2I using SD v1.5, CFG vs CFG++ ($\omega=9.0,\lambda=0.8$). Unnatural depictions of human hands, and incorrect renderings of the text by CFG are corrected in CFG++
 $\omega=2.0,\lambda=0.2$ $\omega=5.0,\lambda=0.4$ $\omega=7.5,\lambda=0.6$ $\omega=9.0,\lambda=0.8$ $\omega=12.5,\lambda=1.0$
@@
 CFG++ (ours) 12.75 0.303 14.95 0.310 17.47 0.312 19.34 0.313 20.88 0.313
-Table 1: Quantitative evaluation (FID, CLIP-similarity) of T2I with SD v1.5
+Table 1: Quantitative evaluation of 50NFE DDIM T2I with SD v1.5 on COCO 10k
 4 Experimental Results
-In this section, we design experiments to show the limitation of CFG and how CFG++ can effectively mitigate these downsides.
-We select Stable Diffusion (rombach2022high) (SD) as our baseline model for its widespread use.
-Experiments are conducted using SD v1.5 from diffusers (patil2022stable) unless specified otherwise.
-Notice that we use the conditioning scale $\lambda\in[0,1]$ for CFG++ whereas for standard CFG, the typical range is $\omega>1.0$.
-When searching for the matched guidance values of $\omega$ and $\lambda$ for a fair comparison, we fix $\lambda=0.2,0.4,0.6,0.8,1.0$ and find the $\omega$ values that produce the images that are of closest proximity in terms of LPIPS distance given the same seed.
-We found that the corresponding values were $\omega=2.0,5.0,7.5,9.0,12.5$, respectively, as shown in Appendix Fig.
-14.
+In this section, we design experiments to show the limitations of CFG and how CFG++ can effectively mitigate these downsides.
+The main experiments were conducted with SD v1.5 or SDXL with 50 NFE DDIM sampling.
+In this regime, we searched for the matching guidance values of $\omega$ and $\lambda$ for a fair comparison.
+We fix $\lambda=0.2,0.4,0.6,0.8,1.0$ and find the $\omega$ values that produce the images that are of closest proximity in terms of LPIPS distance given the same seed.
+We found that the corresponding values were $\omega=2.0,5.0,7.5,9.0,12.5$, respectively.
+Some of the experiments were also conduced with distilled model such as SDXL-turbo, lightning.
 4.1 Text-to-Image Generation
@@
 In Tab.
-1, we observe a constant improvement of the FID metric across all guidance scales, with approximately the same level of CLIP similarity or better.
+1, we report quantitative metrics using 10k images generated from COCO captions (lin2014microsoft).
+Here, we observe a constant improvement of the FID metric across all guidance scales (also see Fig.
+13 for an apples-to-apples comparison), with approximately the same level of CLIP similarity or better.
+Model Metric CFG CFG++ (ours)
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 2: Quant. eval. on accelerated T2I sampling
 The improvements can also be clearly seen in Fig.
@@
 Specifically, we see that unnatural depictions of human hands, and incorrect renderings of the text are corrected in CFG++, a long-standing research question in and of its own (podell2023sdxl; pelykh2024giving; chen2023textdiffuser).
-Figure 8: Interpolating behavior of T2I along various guidance scales using SDXL with CFG++.
-The generated visuals align more significantly with the given textual description as the guidance scale increase to 1.
+We find that the improvement gain from CFG++ is even more dramatic for distilled diffusion models such as SDXL-{turbo, lightning}.
+We quantify the results on the same 5k prompts with 6 NFE sampling.
 In Fig.
-8, we visualize multiple images generated by CFG++ as we increase the guidance scale $\lambda$, illustrating a smooth transition from unconditional to highly conditional sampling.
-Specifically, at higher guidance scales, the target concept is more pronounced, with less emphasis on noisy, unrelated elements.
-This results in images that more closely reflect the specified text conditions.
-For example, in the second row, the images increasingly resemble a "painting by Van Gogh," demonstrating how the generated visuals align more significantly with the given textual description as the guidance scale increase to 1.
+5 and Fig.
+11, we see significant boosts in the quality of the generated images, which is also depicted in the improvements seen in Tab.
+2.
+To show the compatibility of CFG++ with higher-order solvers, we experiment with DPM++2M (lu2022dpmpp) solver with 20 NFE.
+We note that in such low NFE regime, a CFG scale of 5.0 corresponds to the CFG++ scale of 1.0, and we have to slightly extrapolate the value of $\lambda\geq 1.0$ to achieve stronger guidance results.
+In Tab.
+2, we again see that CFG++ yields results that are favorable to standard CFG.
+Figure 5: T2I using SDXL-Lightning, 6 NFE, CFG vs CFG++.
 4.2 Diffusion image Inversion and Editing
-Figure 9: Real image DDIM inversion results using SDv1.5. (a) Reconstructed samples after inversion by CFG and CFG++. (b) Quantitative comparison between CFG and CFG++.
-Figure 10: Real image inversion results at various CFG and CFG++ scales using SDv1.5.
-The image is from COCO dataset.
-We use the text prompt “a white dog is sleeping on a street and a bicycle” during the inversion.
+Figure 6: Inversion and editing results.
+(a) Reconstructed samples after inversion by CFG and CFG++.
+(b) Quantitative comparison between CFG and CFG++ for reconstruction.
+(c) Image editing comparison via SDXL.
 We further explore the effect of CFG++ on DDIM inversion (dhariwal2021diffusion), where the source image is reverted to a latent vector that can reproduce the original image through generative process.
 DDIM inversion is well-known to break down in the usual CFG setting as CFG magnifies the accumulated error of each inversion step (mokady2023null), violating local linearization assumptions (wallace2023edict).
-Remarkably, CFG++ may mitigate this issue.
-This improves inversion and enables better image editing compared to inversion with the original CFG.
+We show that CFG++ mitigates this issue by improving the inversion and image editing capabilities.
 We evaluate our method following the experimental setups in park2024energy and kim2024dreamsampler.
-More details follow.
 Diffusion Image Inversion.
@@
 In Fig.
-9, we illustrate the reconstructed examples (9a) from a real image from COCO data set and computed metrics (9b) for 5k COCO-2014 (lin2014microsoft) validation set.
-Both qualitative and quantitative evaluation demonstrate a consistent improvement on reconstruction performance induced by CFG++.
-Notably, DDIM inversion with CFG++ leads to consistent reconstruction of the source image across all guidance scales, while DDIM inversion with CFG fails to reconstruct it, as shown in Fig.
+6, we illustrate the reconstructed examples (6a) from a real image from COCO data set and computed metrics (6b) for 5k COCO-2014 (lin2014microsoft) validation set.
+Both qualitative and quantitative evaluations demonstrate a consistent improvement on reconstruction performance induced by CFG++.
+Notably, DDIM inversion with CFG++ leads to a consistent reconstruction of the source image across all guidance scales, while DDIM inversion with CFG fails to reconstruct it, as shown in Fig.
 14 for another real image.
@@
 Fig.
-11 compares the image editing result using CFG and CFG++ followed by image inversion.
-In the image editing stage, a word in the source text on the left side is swapped with the target concept, and this modified text is used as condition for sampling.
-For both synthetic and real images, CFG++ successfully edits the target concept while preserving other concepts, such as background.
-In particular, the water splashing or brush in the background, which tend to disappear in the conventional CFG, are maintained through the inversion process by CFG++.
+6c compares the image editing result using CFG and CFG++ followed by image inversion with 5 NFEs.
+In the image editing stage, a word in the source text highlighted by green color is swapped with the target concept, and this modified text is used as the condition for sampling.
+CFG++ successfully edits the target concept while preserving other concepts, such as background.
+In particular, the water splashing in the background, which tends to disappear in the conventional CFG, is maintained through the inversion process by CFG++.
 This emphasizes CFG++’s superior ability to retain specific scene elements that are frequently lost in the standard CFG approach.
-Figure 11: Comparison on image editing via DDIM inversion using SDXL: (top) synthetic image editing, (bottom) real image editing.
+Moreover, standard CFG’s disrupted DDIM inversion leads to saturated and less faithful editing results, whereas CFG++ delivers precise and high-fidelity edits.
 4.3 Text-Conditioned Inverse Problems
@@
 This evaluation utilizes the FFHQ (karras2019style) 512x512 dataset and the text prompt "a high-quality photo of a face".
-Further details about the experimental settings can be found in the Appendix A.
+Further details about the experimental settings can be found in the Appendix C.
 As shown in Tab.
-2, our method mostly outperforms both the vanilla PSLD and PSLD with CFG.
+3, our method mostly outperforms both the vanilla PSLD and PSLD with CFG.
 The superiority is also evident from the Fig.
-12, where CFG++ consistently delivers high-quality reconstructions across all tasks.
+7, where CFG++ consistently delivers high-quality reconstructions across all tasks.
 PSLD with CFG often suffer from artifacts and blurriness.
 Conversely, CFG++ achieves better fidelity, clearly distinguishing between faces and faithfully reproducing fine details like eyelids and hair texture.
-For more results, please refer to the Appendix B.
-Figure 12: Qualitative comparison on various inverse problems using PSLD (rout2024solving) under CFG and CFG++.
+For more results, please refer to the Appendix D.
+Figure 7: Qualitative comparison on various inverse problems using PSLD (rout2024solving) under CFG and CFG++. For more results, please refer to the Appendix D.
 SR (x8) Deblur (motion) Deblur (gauss) Inpaint
@@
 Table 3: Quantitative comparison (FID, LPIPS, PSNR) of PSLD, PSLD with CFG, and PSLD with CFG++ on Latent Diffusion Inverse Solver.
-5 Related Works and Discussion
-Related works.
+5 Related Works and Discussions
 kynkaanniemi2024applying observed that applying CFG guidance at the earlier stages of sampling always led to detrimental effects and drastically reduced the diversity of the samples.
@@
 These findings are orthogonal to ours in that these methods keep the sampling trajectory the same and try to empirically adjust the strength of the guidance, while we aim to design a different trajectory.
-Implications.
-In several studies, attempts at designing a better trajectory of the generative path were made (liu2023flow; lipman2023flow).
-As diffusion models can only estimate the tangent vector of the trajectory at time $t$, it is now a well regarded consensus that minimizing the curvature of the trajectory leads to faster sampling (liu2023flow; lee2023minimizing) and better distillation results (liu2023instaflow).
-In this regard, the excessively high curvature and oscillating behavior of CFG with high guidance scales would hamper the performance of fast sampling and distillation, which was constantly reported in the literature as difficulties in extending distillation methods (salimans2022progressive; song2023consistency) to the latent space (meng2023distillation; luo2023latent; chen2024pixart).
-Straightening the diffusion path prior to distillation with Reflow (liu2023flow) exhibited much better results than aiming for direct distillation.
-In this regard, CFG++ can also be considered a better alternative to CFG for enhanced distillation performance.
+More recently, bradley2024classifier observed that the CFG score is not a valid denoising direction, showing that in the asymptotic limit, CFG sampling can be considered as a specific type of predictor-corrector (PC) sampling, dubbed PCG.
+PCG shares a similar spirit with CFG++ in that for the mainstream of sampling, one defers from the use of mixed CFG score (we use unconditional score, PCG uses conditional score) to avoid invalid directions.
+While PCG derives an interesting connection of PC sampling with CFG, the proposed algorithm does not improve the sampling performance.
+In contrast, CFG++ improves the sample quality without additional computation overhead.
+Since CFG++ attempts to make adjustments of CFG using a linear combination of score functions, it is not surprising to see that CFG++ sampling can be achieved by setting a time varying schedule.
+Specifically, by setting a time-varying schedule $\omega_{t}$ as $\omega_{t}=-\gamma_{t}/\xi_{t}$, where $\gamma_{t}=\frac{\sqrt{\bar{\alpha}_{t-1}}\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}},\xi_{t}=\sqrt{1-\bar{\alpha}_{t-1}}-\frac{\sqrt{\bar{\alpha}_{t-1}}\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}$, one can achieve the same effect with CFG as CFG++.
+However, this specific choice of time-varying guidance scale has not been reported in the literature, as it is a schedule that is relatively complex to be drawn heuristically.
+We further show in Appendix B that such choice removes undesirable oscillatory behavior in the evolution of the posterior mean.
 6 Conclusion
@@
 $\lambda\in[0.0,1.0]$, might suffice successful guidance with a proper geometric correction.
-Observing that the original CFG suffers from off-manifold issues during sampling, we propose a simple but surprisingly effective fix: using ${\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})$ instead of ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$ in updating ${\boldsymbol{x}}_{t-1}$.
-This change translates the conditional guidance from extrapolating to interpolating between unconditional and conditional sampling trajectories, leading to more interpretable guidance in contrast to the conventional CFG which heavily relies on heuristic $\omega$ scaling.
+Observing that the original CFG suffers from off-manifold issues during sampling, we propose a simple but surprisingly effective fix.
+This change translates the conditional guidance from extrapolating to interpolating between unconditional and conditional sampling trajectories, leading to more interpretable guidance in contrast to the conventional CFG which relies on heuristic $\omega$ scaling.
 Given that CFG++ mitigates off-manifold issues, it may be beneficial for other downstream applications requiring accurate latent denoised estimate representations, e.g.
 DIS, as demonstrated in our experiments.
-Appendix A Experimental details
-A.1 Text-Conditioned Inverse Problems
+Acknowledgments
+We thank Jinho Chang for the helpful discussions.
+Appendix A Extension of CFG++ to other solvers
+In this section, we discuss ways in which we can apply CFG++ to a more diverse set of ODE/SDE solvers.
+We consider solving the variance exploding (VE) PF-ODE as presented in (1), as re-parametrization VP diffusion models can easily recover the VE formulation, as often implemented in widely used frameworks such as https://github.com/crowsonkb/k-diffusion.
+Following the notation in lu2022dpmpp, we consider a sequence of timesteps $\{t_{i}\}_{i=0}^{M}$, where $t_{0}=T$ denotes the initial starting point of the reverse sampling (i.e.
+Gaussian noise).
+Euler (karras2022elucidating)
+The construction is the same as in DDIM. We include it here for completeness. The update step reads
+${\boldsymbol{x}}_{t_{i+1}}$ $={\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t_{i}})+\frac{{\boldsymbol{x}}_{t_{i}}-{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t_{i}})}{\sigma_{t_{i}}}\cdot\sigma_{t_{i+1}},\quad$ (CFG)
+${\boldsymbol{x}}_{t_{i+1}}$ $={\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t_{i}})+\frac{{\boldsymbol{x}}_{t_{i}}-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t_{i}})}{\sigma_{t_{i}}}\cdot\sigma_{t_{i+1}},\quad$ (CFG++)
+Euler Ancestral
+Euler Ancestral sampler follows Euler sampler, but introduces stochasticity by taking larger steps and then adding a slight amount of noise. Adjusting with CFG++ is straightforward.
+${\boldsymbol{x}}_{t_{i+1}}$ $={\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t_{i}})+\frac{{\boldsymbol{x}}_{t_{i}}-{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t_{i}})}{\sigma_{t_{i}}}\cdot(\sigma_{t_{d_{i}}}-\sigma_{t_{i}})+\sigma_{t_{i}}{\boldsymbol{\epsilon}},\quad$ (CFG)
+${\boldsymbol{x}}_{t_{i+1}}$ $={\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t_{i}})+\frac{{\boldsymbol{x}}_{t_{i}}-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t_{i}})}{\sigma_{t_{i}}}\cdot(\sigma_{t_{d_{i}}}-\sigma_{t_{i}})+\sigma_{t_{i}}{\boldsymbol{\epsilon}},\quad$ (CFG++)
+where $t_{i}>t_{d_{i}}>t_{i+1}$ and ${\boldsymbol{\epsilon}}\sim{\mathcal{N}}(0,{\bm{I}})$.
+DPM-solver++ 2M (lu2022dpmpp)
+Define $\sigma_{t}:=e^{-t},\,h_{i}:=t_{i}-t_{i-1},$ and $r_{i}:=h_{i-1}/h_{i}$. After initializing ${\boldsymbol{x}}_{t_{0}}$ with Gaussian noise, the first iteration is given by
+${\boldsymbol{x}}_{t_{1}}$ $={\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t_{0}})+e^{-h_{1}}({\boldsymbol{x}}_{t_{0}}-{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t_{0}}))\quad$ (CFG)
+${\boldsymbol{x}}_{t_{1}}$ $={\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t_{0}})+e^{-h_{1}}({\boldsymbol{x}}_{t_{0}}-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t_{0}}))\quad$ (CFG++)
+The following iterations by using the standard CFG reads
+$\bm{D}_{i}$ $={\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t_{i-1}})+\frac{1}{2r_{i}}\left({\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t_{i-1}})-{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t_{i-2}})\right),$ (23)
+${\boldsymbol{x}}_{t_{i}}$ $=e^{-h_{i}}{\boldsymbol{x}}_{t_{i-1}}-(e^{-h_{i}}-1)\bm{D}_{i}.$ (24)
+Rearranging (23), (24), we can rewrite the update steps as
+${\boldsymbol{x}}_{t_{i}}$ $={\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t_{i-1}})-e^{-h_{i}}{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t_{i-1}})+\frac{1-e^{-h_{i}}}{2r_{i}}\left({\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t_{i-1}})-{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t_{i-2}})\right)+e^{-h_{i}}{\boldsymbol{x}}_{t_{i-1}}.$ (25)
+Notice that in order to apply CFG++ to (25), we should only keep the first term as the conditional Tweedie, and use the unconditional estimates for the rest of the components. i.e.
+${\boldsymbol{x}}_{t_{i}}$ $={\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t_{i-1}})-e^{-h_{i}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t_{i-1}})+\frac{1-e^{-h_{i}}}{2r_{i}}\left({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t_{i-1}})-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t_{i-2}})\right)+e^{-h_{i}}{\boldsymbol{x}}_{t_{i-1}}$ (26)
+DPM-solver++ 2S (lu2022dpmpp)
+With the same choices of $\sigma_{t},h_{i}$, we additionally define the timesteps $\{s_{i}\}_{i=1}^{M}$ with $t_{i}>s_{i+1}>t_{i+1}$.
+Further, let $r_{i}=\frac{s_{i}-t_{i-1}}{t_{i}-t_{i-1}}$.
+Using standard CFG, the iteration reads
+$\bm{u}_{i}$ $=e^{-r_{i}h_{i}}{\boldsymbol{x}}_{t_{i-1}}+(1-e^{-r_{i}h_{i}}){\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t_{i-1}})$ (27)
+${\boldsymbol{x}}_{t_{i}}$ $={\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t_{i-1}})-e^{-h_{i}}{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t_{i-1}})+\frac{1-e^{-h_{i}}}{2r_{i}}\left({\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}(\bm{u}_{i})-{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t_{i-1}})\right)+e^{-h_{i}}{\boldsymbol{x}}_{t_{i-1}}$ (28)
+Applying the general transition rule from CFG to CFG++ as introduced in (15), we can keep all of the Tweedie estimates to be unconditional, and only change the first term of (28), i.e.
+$\bm{u}_{i}$ $=e^{-r_{i}h_{i}}{\boldsymbol{x}}_{t_{i-1}}+(1-e^{-r_{i}h_{i}}){\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t_{i-1}})$ (29)
+${\boldsymbol{x}}_{t_{i}}$ $={\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t_{i-1}})-e^{-h_{i}}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t_{i-1}})+\frac{1-e^{-h_{i}}}{2r_{i}}\left({\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}(\bm{u}_{i})-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t_{i-1}})\right)+e^{-h_{i}}{\boldsymbol{x}}_{t_{i-1}}.$ (30)
+For the ancestral version of DPM-solver++ 2S, we can follow the general transition rule, as was shown for Euler $\rightarrow$ Euler Ancestral.
+Appendix B Evolution of the posterior mean: CFG vs. CFG++
+Recall that one can equivalently view the evolution of the posterior mean ${\mathbb{E}}[{\boldsymbol{x}}_{0}|{\boldsymbol{x}}_{t}]$ rather than the noisy variables ${\boldsymbol{x}}_{t}$.
+In this context, we derive the sequential evolution of conditional posterior mean ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}$ and ${\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}$ through time $t$ to further understand the underlying behavior of the proposed sampling.
+Proposition 1.
+Let $d{\boldsymbol{z}}({\boldsymbol{x}}_{t}):={\boldsymbol{z}}({\boldsymbol{x}}_{t})-{\boldsymbol{z}}({\boldsymbol{x}}_{t+1})$ denote the discrete time evolution of some random variable ${\boldsymbol{z}}$ at time $t$.
+Then, the evolution of ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}$ of CFG and ${\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}$ of CFG++ is given by
+$d{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$ $=\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}d{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})+\omega(\Delta({\boldsymbol{x}}_{t},{\boldsymbol{c}})-\Delta({\boldsymbol{x}}_{t+1},{\boldsymbol{c}}))$ (31)
+$d{\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})$ $=\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}\underbrace{d{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})}_{\text{uncond.
+shift}}+\lambda\underbrace{\Delta({\boldsymbol{x}}_{t},{\boldsymbol{c}})}_{\text{cond.
+shift}},$ (32)
+where $\Delta({\boldsymbol{x}}_{t},{\boldsymbol{c}}):={\hat{\boldsymbol{x}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})$.
+Proof.
+We start by writing the iteration from $t+1\rightarrow t$
+${\boldsymbol{x}}_{t}=\sqrt{\bar{\alpha}_{t}}{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t+1})+\sqrt{1+\bar{\alpha}_{t}}{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t+1}).$ (33)
+The Tweedie estimate for the next step is then written as
+${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$ $=\frac{{\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})}{\sqrt{\bar{\alpha}_{t}}}$ (34)
+$=\frac{\sqrt{\bar{\alpha}_{t}}{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t+1})+\sqrt{1-\bar{\alpha}_{t}}({\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t+1})-{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t}))}{\sqrt{\bar{\alpha}_{t}}}$ (35)
+$={\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t+1})+\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}\Bigl{[}{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t+1})-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})+\omega({\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t+1})-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t+1}))$
+$-\omega({\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t}))\Bigr{]}.$ (36)
+Using the relation ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})=-(\sqrt{\bar{\alpha}_{t}}{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})-{\boldsymbol{x}}_{t})/\sqrt{1-\bar{\alpha}_{t}}$, we have
+$d{\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}({\boldsymbol{x}}_{t})$ $=\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}d{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})+\omega\left({\hat{\boldsymbol{x}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})\right)-\omega\left({\hat{\boldsymbol{x}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t+1})-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t+1})\right).$ (37)
+Similarly, for CFG++,
+${\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})$ $=\frac{{\boldsymbol{x}}_{t}-\sqrt{1-\bar{\alpha}_{t}}{\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})}{\sqrt{\bar{\alpha}_{t}}}$ (38)
+$=\frac{\sqrt{\bar{\alpha}_{t}}{\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t+1})+\sqrt{1-\bar{\alpha}_{t}}({\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t+1})-{\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t}))}{\sqrt{\bar{\alpha}_{t}}}$ (39)
+$={\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t+1})+\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}\left[{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t+1})-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})-\lambda({\hat{\boldsymbol{\epsilon}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t}))\right].$ (40)
+Hence,
+$d{\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}({\boldsymbol{x}}_{t})$ $=\frac{\sqrt{1-\bar{\alpha}_{t}}}{\sqrt{\bar{\alpha}_{t}}}d{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})+\lambda({\hat{\boldsymbol{x}}_{\boldsymbol{c}}}({\boldsymbol{x}}_{t})-{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{x}}_{\varnothing}}({\boldsymbol{x}}_{t})),$ (41)
+Proposition 1 implies that the CFG++ update of ${\color[rgb]{0.165,0.616,0.561}\definecolor[named]{pgfstrokecolor}{rgb}{0.165,0.616,0.561}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\lambda}}$ is decomposed into two shift terms: 1) $d{\color[rgb]{0.208,0.565,0.953}\definecolor[named]{pgfstrokecolor}{rgb}{0.208,0.565,0.953}\hat{\boldsymbol{\epsilon}}_{\varnothing}}({\boldsymbol{x}}_{t})$ represents the difference between consecutive unconditional scores (i.e.
+unconditional shift), and 2) $\Delta({\boldsymbol{x}}_{t},{\boldsymbol{c}})$ denotes the direction of conditional guidance at time $t$ (i.e.
+conditional shift).
+The conditional shift term is multiplied by a small interpolation factor $\lambda$ in the case of CFG++, inducing a small nudge toward the condition.
+CFG ${\color[rgb]{0.906,0.435,0.318}\definecolor[named]{pgfstrokecolor}{rgb}{0.906,0.435,0.318}\hat{\boldsymbol{x}}_{\boldsymbol{c}}^{\omega}}$, on the other hand, has the same unconditional shift, but has an oscillatory behavior for the conditional shift.
+The difference between CFG/CFG++ sampling arises from the unexpected additional shift from the previous timestep $t+1$ that exists for the CFG decomposition: $-\Delta({\boldsymbol{x}}_{t+1},{\boldsymbol{c}})$, and the scaling constant $\omega$.The initial conditional shift $\omega\Delta({\boldsymbol{x}}_{t},{\boldsymbol{c}})$ with a large $\omega$ pushes the trajectory off the manifold, but cancels some of its effects by subtracting the conditional shift from the previous step $\omega\Delta({\boldsymbol{x}}_{t+1},{\boldsymbol{c}})$.
+The compounded vector $\Delta({\boldsymbol{x}}_{t},{\boldsymbol{c}})-\Delta({\boldsymbol{x}}_{t+1},{\boldsymbol{c}})$ does induce a nudge closer to the condition but requires a large value of $\omega$ to have a meaningful effect, and thus is hard to interpret.
+Appendix C Experimental details
+C.1 Text-Conditioned Inverse Problems
 Problem settings.
@@
 Table 4: Hyperparameters for PSLD (rout2024solving) with CFG++ and corresponding CFG scale $\omega$.
-Appendix B Further experimental results
-B.1 T2I
+Appendix D Further experimental results
+D.1 T2I
 Since the range of guidance scales for CFG and CFG++ is different, we matched the guidance values of $\omega$ and $\lambda$ for comparison by computing the LPIPS distance using the same seed.
-Fig.
-14 displays an example of generated images given the same text prompt and seed, while varying the guidance scale.
 Fig.
@@
 However, the CFG++ addresses this issue by adjusting the DDIM re-noising step and the guidance scale.
+Figure 9: Enhanced T2I results by SDXL ($\omega=9.0,\lambda=0.8$) with CFG++.
+Under CFG, the lion cub is not visible (top-left), the dog appears with two tails (top-right), the goggles have an unusual shape (bottom-left), and the tree trunk is folded (bottom-right).
+These artifacts are absent in those produced by CFG++.
+Figure 10: T2I using SD v1.5, CFG vs CFG++ ($\omega=9.0,\lambda=0.8$). Unnatural depictions of human hands, and incorrect renderings of the text by CFG are corrected in CFG++.
+Figure 11: T2I using SDXL-Lightning, 6 NFE, CFG vs CFG++. The overall image quality and sophistication have improved with CFG++.
 Figure 12: The discrete evolution of the posterior mean in CFG and CFG++. Denoised estimates in latent space are decoded into pixel space at each timestep using SDXL.
-Figure 14: Example of generated images with various guidance scales by CFG and CFG++. The respective guidance scales are chosen to best match the generated images by CFG and CFG++.
-B.2 Real Image Editing
+Figure 13: Interpolating behavior of T2I along various guidance scales using SD v1.5 with CFG and CFG++
+Figure 14: Real image inversion results at various CFG and CFG++ scales using SDv1.5.
+The image is from COCO dataset.
+We use the text prompt “a white dog is sleeping on a street and a bicycle” during the inversion.
+D.2 Real Image Editing
 In Fig.
@@
 Figure 17: Comparison on real image editing via DDIM inversion under CFG and CFG++ using SDXL. Cat $\rightarrow$ Cat with glasses.
-B.3 Text-Conditioned Inverse Problems
+D.3 Text-Conditioned Inverse Problems
 In Fig.